### v3

Abstract
Entry-exit dynamics are crucial in modeling crowd movement. Here, we present a novel first-order, stationary mean-field game (MFG) model on bounded domains that accurately captures entry-exit dynamics. In our model, the interior dynamics are governed by a standard first-order stationary MFG system: a first-order Hamilton-Jacobi equation coupled with a transport equation. The model incorporates mixed boundary conditions that correspond to an entry region $\Gamma_{N}$ and an exit region $\Gamma_{D}$. A Neumann condition on $\Gamma_{N}$ prescribes the agent inflow via a non-homogeneous flux term, $j(x)$; a no-entry condition on $\Gamma_{D}$ restricts this boundary region to exit only, preventing inward flow; finally, in $\Gamma_{D}$, we prescribe an upper bound on the exit cost combined with a complementary contact-set condition.
This contact-set condition identifies boundary points where the value function attains the exit cost (contact points) versus points where the no-penetration condition prevents artificial inflows (non-contact points). However, as our examples show, contact does not necessarily imply that exit occurs. This mixed approach overcomes the limitations of classical Dirichlet conditions, which can artificially force boundary points to act as both entry and exit sites.
We analyze the system using a variational formulation, applying the direct method of calculus of variations to establish the existence of solutions under minimal regularity assumptions. Furthermore, we prove the uniqueness of the gradient of the value function (particularly in regions with positive agent density) and the uniqueness of the density function.
Several examples, including cases in one and two dimensions, illustrate first-order MFG phenomena such as the formation of empty regions (where agent density vanishes) and the proper assignment of entry and exit roles. These results establish a rigorous mathematical foundation for modeling realistic entry-exit scenarios.
1 Introduction
This paper studies first-order stationary Mean-Field Games (MFGs) in bounded domains under nonstandard mixed boundary conditions (BCs). These boundary conditions induce a novel MFG structure that resolves an inherent issue with standard Dirichlet conditions: a boundary point may inadvertently serve as an entrance instead of an exit. This MFG model a scenario where agents can enter or exit only through designated boundary portions. This new MFG system is relevant in various applications, such as crowd dynamics, where individuals may join or leave a venue through different gates, or economic models, where firms may enter or exit a market under certain conditions. One of the paper’s main goals is to show that the mixed BCs lead to a well-defined MFG formulation. The precise problem we consider is as follows.
Problem 1 (The Mean-Field Game (MFG)).
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded, open, and connected $C^{1}$–domain whose boundary $\Gamma=\partial\Omega$ is partitioned into disjoint sets $\Gamma_{D}$ and $\Gamma_{N}$; that is, $\Gamma=\mkern 1.0mu\overline{\mkern-1.0mu\Gamma\mkern-1.0mu}\mkern 1.0mu_{D}\cup\mkern 1.0mu\overline{\mkern-1.0mu\Gamma\mkern-1.0mu}\mkern 1.0mu_{N}$. Let
$j\colon\Gamma_{N}\to[0,\infty),\enskip\,\psi\colon\Gamma_{D}\to\mathbb{R},\enskip\,H\colon\Omega\times\mathbb{R}^{d}\to\mathbb{R},\enskip\,\text{and }\enskip g\colon[0,\infty)\to\mathbb{R}$
be given functions such that the $C^{1}$–mapping $p\mapsto H(x,p)$ is convex and that $g$ is continuous and increasing. Solve the following system for $(m,u)$.
$\begin{cases}H(x,Du)=g(m)&\text{ in }\Omega,\\ -\operatorname*{div}(mD_{p}H(x,Du))=0&\text{ in }\Omega,\end{cases}$ (1.1)
with the boundary conditions
$\begin{cases}mD_{p}H(x,Du)\cdot\nu=j(x)&\text{ on }\Gamma_{N},\\ mD_{p}H(x,Du)\cdot\nu\leq 0\,\text{ and }\,u(x)\leq\psi(x)&\text{ on }\Gamma_{D},\\ (\psi-u)mD_{p}H(x,Du)\cdot\nu=0&\text{ on }\Gamma_{D}.\end{cases}$ (1.2)
The previous problem corresponds to the equilibrium state of a game in a bounded domain with a boundary divided into two regions: an entry region $\Gamma_{N}$ and an exit region $\Gamma_{D}$. Here, $\Gamma_{D}$ and $\Gamma_{N}$ are disjoint, smooth $(d-1)$–dimensional manifolds, relatively open with respect to $\Gamma$, with
$\left|\mkern 1.0mu\overline{\mkern-1.0mu\Gamma\mkern-1.0mu}\mkern 1.0mu_{D}\cap\mkern 1.0mu\overline{\mkern-1.0mu\Gamma\mkern-1.0mu}\mkern 1.0mu_{N}\right|=0.$
On $\Gamma_{N}$, we prescribe the inflow of agents via a Neumann condition. The non-homogeneous Neumann condition (i.e., $j(x)\not\equiv 0$ on $\Gamma_{N}$) is essential to ensure a nontrivial influx of agents, preventing the trivial solution $m\equiv 0$. On $\Gamma_{D}$, we enforce a relaxed Dirichlet condition on the value function; the complementary condition on $\Gamma_{D}$ ensures it functions solely as an exit region with $\psi$ as exit cost, avoiding artificial inflows. We denote by
$\Gamma_{D}^{1}:=\{x\in\Gamma_{D}:u(x)=\psi(x)\}$
the contact set (where the value function attains the exit cost) and by
$\Gamma_{D}^{0}:=\Gamma_{D}\setminus\Gamma_{D}^{1}$
the non-contact set. In the non-contact set, the condition $(\psi-u)mD_{p}H(x,Du)\cdot\nu=0$ implies zero flow $mD_{p}H(x,Du)\cdot\nu=0$, thus preventing artificial inflows through the exit that can arise if we impose a strict Dirichlet boundary condition. Additional assumptions on functions $H$, $g$, $j$ and $\psi$ are given in Section 2. In Section 3.1, we present several examples that illustrate key phenomena predicted by our model, such as the formation of empty regions.
1.1 The derivation of the model
Here, we show how Problem 1, specifically the PDE system (1.1)-(1.2), emerges as the stationary state of a time-dependent coupled optimal control and transport problem (see [GPV16]). For simplicity, we focus on the special case where:
$H(x,p)=\frac{1}{2}|p|^{2},\quad g(m)=m.$
The corresponding Lagrangian is:
$L(v)=\sup_{p}\{-p\cdot v-\frac{1}{2}|p|^{2}\}=\frac{1}{2}|v|^{2}.$
We consider the setting of Problem 1, namely a domain $\Omega\subset\mathbb{R}^{d}$ with boundary $\Gamma=\overline{\Gamma}_{D}\cup\overline{\Gamma}_{N}$, let $j\geq 0$ be a $C^{1}$ function defined on $\Gamma_{N}$, and $\psi$ a continuous function defined on $\Gamma_{D}$.
1.1.1 The Optimal Control
Consider a population whose density $m(t,x)$ is known for all time $t>0$, and consider a typical player in this population whose trajectory $\mathbf{x}$ is governed by the following dynamics:
$\begin{cases}\dot{\mathbf{x}}(\tau)=\mathbf{v}(\tau,\mathbf{x}(\tau))&\text{for }t<\tau<T_{E}\\ \mathbf{x}(t)=x&\mathbf{x}(T_{E})\in\Gamma_{D}.\end{cases}$ (1.3)
The effect of the population on the player is two-fold. First, the player is allowed to choose $\mathbf{v}$ only from the admissible set
$\mathcal{V}=\{\mathbf{v}\in C^{0}([t,+\infty)\times\mkern 1.0mu\overline{\mkern-1.0mu\Omega\mkern-1.0mu}\mkern 1.0mu,\mathbb{R}^{d})\colon-m(t,x)\mathbf{v}(t,x)\cdot\nu=j(x)\text{ on }\Gamma_{N},\enskip m\mathbf{v}\cdot\nu\geq 0\text{ on }\Gamma_{D}\},$
which enforces the incoming flow condition $j$ and prevents (normal) entry through $\Gamma_{D}$. Second, the player optimizes the control according to the following optimal control problem.
$\underset{\mathbf{v}\in\mathcal{V}}{\text{Minimize}}\quad\mathcal{J}[\mathbf{v};t,x]:=\int_{t}^{T_{E}}\frac{1}{2}|\mathbf{v}(\tau,\mathbf{x{(\tau)})}|^{2}+m(\tau,\mathbf{x{(\tau)}}){\,\,\rm d}\tau+\psi(\mathbf{x}(T_{E})),$ (1.4)
where the exit time $T_{E}$ is chosen by the player such that $\mathbf{x}(T_{E})\in\Gamma_{D}$. Under enough smoothness assumptions, the classical theory of optimal control ensures that the value function
$u(t,x):=\min_{\mathbf{v}}\mathcal{J}[\mathbf{v};t,x],$
solves the Hamilton-Jacobi equation
$-\partial_{t}u(t,x)+\frac{1}{2}|Du(t,x)|^{2}-m(t,x)=0\quad\text{ for }t\in(0,\infty),\,x\in\Omega.$ (1.5)
Furthermore, it also gives us the optimal velocity field in the feedback form
$\mathbf{v}^{*}:=-Du\in\mathcal{V},$ (1.6)
In the optimal control problem (1.4), the trajectories $\mathbf{x}(t)$ evolve within $\bar{\Omega}$ according to (1.3). When a player reaches the boundary in a transversal direction—i.e., the velocity has a strictly positive normal component— the player is forced to exit immediately. In this case, $\mathbf{v}(T_{E},\mathbf{x}(T_{E}))\cdot\nu>0$ and the player must incur the exit cost $u(T_{E},\mathbf{x}(T_{E}))={\psi(\mathbf{x}(T_{E}))}$. If a player reaches the boundary tangentially, they are not compelled to exit immediately; instead, they may continue along the boundary—potentially traversing parts of $\Gamma_{N}$ where $j=0$, and choose a later optimal exit time, $T_{E}$. Because tangential motion does not mandate an immediate exit, the boundary conditions must not force the value function to exactly match the exit cost, rendering a strict Dirichlet boundary condition on $\Gamma_{D}$ unsuitable. Instead, we impose the relaxed boundary condition
$u(t,x)\;\leq\;\psi(x),\quad x\,\in\,\Gamma_{D},$
thereby accounting for the possibility of the absence of an enforced exit at those points.
Thus, $u$ solves (1.5) with the boundary conditions:
$\begin{cases}m(t,x)Du(t,x)\cdot\nu(x)=j(x)&\text{ for }t\in(0,\infty),\,x\in\Gamma_{N}\\ u(t,x)\leq{\psi(x)},\quad{m(t,x)}Du(t,x)\cdot\nu(x)\leq 0&\text{ for }t\in(0,\infty),\,x\in\Gamma_{D}\\ {(\psi(x)-u(t,x))}{m(t,x)}Du(t,x)\cdot\nu(x)=0&\text{ for }t\in(0,\infty),\,x\in\Gamma_{D}\end{cases}$ (1.7)
Here, the normal derivative is weighted by the density $m(t,x)$ to ensure that the flux condition is enforced only at boundary points where agents are present. In particular, if $m(t,x)=0$ at a point on the boundary, the normal velocity is immaterial, and the flux condition is automatically satisfied.
1.1.2 Transport
The second part of the modeling concerns the collective dynamics of the population. If all agents act optimally, their trajectories will follow the optimal velocity field (1.6). Thus, given an initial population density $m(0,x)=m_{0}$, the flow of the vector field $-Du(t,x)$ transports this density according to
$m(t,\cdot)=\Phi(t,\cdot)_{\#}m_{0},$
where $\Phi$ is the flow field of the ODE
$\dot{\mathbf{x}}(t)=-Du(t,\mathbf{x}(t))\qquad\text{for }t>0,$
and ##\## denotes the pushforward. More explicitly, let $\Phi(t,\cdot)^{-1}$ be the preimage of $\Phi$ as a function of $x$ and constant $t$. Then, for any $\varphi\in C_{c}^{\infty}(\Omega)$, and any $t$ for which $A=\Phi(t,\cdot)^{-1}(\operatorname{supp}(\varphi))$ is well defined,
$\int_{\Omega}\varphi(x)\,m(t,x){\,\,\rm d}x=\int_{A}\varphi(\Phi(t,y))\,m_{0}(y){\,\,\rm d}y,$
Taking into account that $\varphi(\Phi(t,y))$ vanishes in $\partial A$, differentiating with respect to $t$, we obtain
$\begin{split}\int_{\Omega}\varphi(x)\,\partial_{t}m(t,x){\,\,\rm d}x&=\int_{A}\partial_{t}\Phi(t,y)\cdot D\varphi(\Phi(t,y))\,m_{0}(y){\,\,\rm d}y\\ &=-\int_{A}Du(t,\Phi(t,y))\cdot D\varphi(\Phi(t,y))\,m_{0}(y){\,\,\rm d}y.\end{split}$
Applying the definition of $m(t,\cdot)=\Phi(t,\cdot)_{\#}m_{0}$ again, we obtain
$\int_{\Omega}\varphi(x)\,\partial_{t}m(t,x){\,\,\rm d}x=-\int_{\Omega}Du(t,x)\cdot D\varphi(x)\,m(t,x){\,\,\rm d}x,$
for all $\varphi\in C_{c}^{\infty}(\Omega)$. This identity is the weak formulation of the transport equation:
$\partial_{t}m(x,t)-\operatorname*{div}(Du(t,x)\,m(t,x))=0.$ (1.8)
1.1.3 Stationary State
In the stationary case, the value function $u(x)$ and the density $m(x)$ become time-independent, and the Hamilton–Jacobi equation (1.5) reduces to
$\frac{1}{2}|Du(x)|^{2}-m(x)=0,$
while the transport equation (1.8) becomes
$-\operatorname*{div}\big{(}m(x)Du(x)\big{)}=0.$
Together with the boundary conditions (1.7), these equations are the stationary mean field game system described in Problem 1.
1.2 Background and prior work
The MFG model (1.1)–(1.2) belongs to a broad class of differential games that describe the distribution of a large population of competing rational agents. Early works, such as [Aum64], [JR88], and [BB92], investigated various aspects of these games; however, these approaches initially received limited attention from the PDE community (see [CHM18]). In 2006–2007, Lasry and Lions introduced a new formulation that incorporated PDE tools and optimal control theory [LL06a, LL06b, LL07], and independently, Huang et al. provided complementary insights from a control theory perspective [HMC06, HMC07]. These seminal contributions laid the foundation for modern MFG theory, where each player seeks to minimize a cost functional based on statistical information about the others.
In these games, the resulting models have two main unknowns: the value function $u$, representing the minimum cost achievable by a typical agent, and the agents’ distribution in the game domain, denoted by $m$. As in our model, (1.1), these two quantities are usually determined by two coupled PDEs: a Hamilton–Jacobi equation and a transport or Fokker-Planck equation. The Hamilton–Jacobi equation characterizes the value function for a typical agent. Usually, the Hamilton-Jacobi equation is either a first or second-order PDE. First-order models, like the one we consider here, arise from deterministic scenarios where a player’s path depends solely on their control and initial state. The first-order problem with local coupling in the deterministic limit is studied in [CG15], where the authors establish the existence and uniqueness of weak solutions under general structure conditions and with periodic boundary conditions (also see [GM18] for further regularity results). Like the one in the current paper, their approach is variational. Second-order models, which include a Laplacian or elliptic term, stem from stochastic models where a player’s state is a random process influenced by Brownian noise (for more details and additional references, see, for example, [GPV16]). Stationary games can be interpreted as the equilibrium state of a long-running game or can arise from ergodic models focusing on the long-term average of a cost functional, as discussed in [LL06a].
In [LL06a, LL06b], MFGs were studied under periodic boundary conditions. Subsequent investigations have focused on models with either Neumann or Dirichlet boundary conditions motivated by practical applications. For instance, using monotonicity methods, [FGT19] addresses the first-order Dirichlet problem, proving the existence of weak solutions. However, because these formulations lack prescribed Neumann conditions, the boundary does not dictate a specific inflow or outflow, and any segment may function as an entrance or exit. Similarly, in [BS24] and [Ber17], the authors examine Dirichlet problems in the contexts of pedestrian evacuation and debt refinancing models, and optimal stopping, respectively. For second-order problems, Neumann conditions have been treated in [Cir15] and [MS18], the latter employing a variational approach that accommodates non-monotone local coupling functions. Related studies on density constraints (e.g., [San12], [CMS16]) also use variational techniques to derive MFGs from optimality conditions and establish the existence of solutions.
While prior research has separately addressed MFGs with either first-order Hamilton–Jacobi equations or isolated Dirichlet/Neumann boundary conditions, no study has yet formulated a model that simultaneously incorporates prescribed Neumann inflow and Dirichlet exit constraints. Our work fills this gap by developing a rigorous variational framework that guarantees the existence (and partial uniqueness) of solutions under explicit structural assumptions.
1.3 Main results
This paper demonstrates that incorporating nonstandard mixed boundary conditions, assigning one part of the boundary for agent entry (Neumann conditions) and another for agent exit (relaxed Dirichlet conditions), results in a well-defined stationary MFG formulation. Under minimal regularity assumptions, we introduce a notion of weak solutions (Definition 2.11) consistent with prior literature and accommodating these mixed boundary conditions.
Building upon this concept, we develop a variational formulation that naturally incorporates the desired boundary behaviors. This approach allows us to:
  • Establish the existence of solutions;
  • Prove a partial uniqueness result (uniqueness of the gradient where the density is positive);
  • Identify the minimal regularity conditions required for well-posedness and, in particular, provide a meaning for the Neumann boundary conditions.
In Section 2, we detail our technical assumptions, define our weak solution framework, and outline formal PDE estimates that motivate our choice of functional spaces. A critical technical challenge is rigorously interpreting the Neumann boundary conditions, as $mD_{p}H(x,Du)$ is only guaranteed to be in a Lebegue space. However, because it is divergence-free, we can interpret the Neumann boundary conditions through the normal trace (see Theorem 2.8).
To illustrate the limitations of standard Dirichlet conditions, Section 3 discusses concrete examples. Moreover, additional examples illustrate phenomena such as the formation of regions with vanishing density (see Section 3.3).
We now introduce our variational formulation, which naturally yields the appropriate boundary conditions:
Problem 2 (The variational formulation).
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded, open, and connected $C^{1}$–domain whose boundary $\Gamma=\partial\Omega$ is partitioned into disjoint sets $\Gamma_{D}$ and $\Gamma_{N}$; that is, $\Gamma=\mkern 1.0mu\overline{\mkern-1.0mu\Gamma\mkern-1.0mu}\mkern 1.0mu_{D}\cup\mkern 1.0mu\overline{\mkern-1.0mu\Gamma\mkern-1.0mu}\mkern 1.0mu_{N}$. Let
$j\colon\Gamma_{N}\to[0,\infty),\quad\psi\colon\Gamma_{D}\to\mathbb{R},\quad H\colon\Omega\times\mathbb{R}^{d}\to\mathbb{R},\quad\text{and}\quad G\colon\mathbb{R}\to\mathbb{R}$
be given functions such that the $C^{1}$–mapping $p\mapsto H(x,p)$ is convex and that $G$ is $C^{1}$, convex, and increasing. Minimize the functional
$\mathcal{I}[w]:=\int_{\Omega}G\left(H(x,Dw)\right){\,\,\rm d}x-\int_{\Gamma_{N}}jw{\,\,\rm d}s,$ (1.9)
over the admissible set
$\mathcal{U}=\big{\{}w\in W^{1,\gamma}(\Omega):\,w(x)\leq\psi(x)\text{ for }x\in\Gamma_{D}\big{\}}.$
Our main result establishes that solutions to the MFG system (Problem 1) exist and that they correspond precisely to the minimizers of Problem 2. To achieve this, in Section 4, we prove the following existence theorem:
Theorem 1.1.
Under Assumptions 1-4 and Assumption 5’, there exists $u\in\mathcal{U}$ such that
$\begin{split}\mathcal{I}[u]=\inf_{w\in\mathcal{U}}\mathcal{I}[w].\end{split}$
The proof employs the direct method of calculus of variations, addressing the key challenge of establishing coercivity under mixed boundary conditions. Moreover, this variational framework provides a justification of the regularity estimates in Section 2.4 and is crucial for the existence of the normal trace for the flux, thus giving precise meaning to the Neumann boundary condition (see Theorem 2.8).
Furthermore, we have the following correspondence theorem:
Theorem 1.2.
Under the Assumptions 1-5 and Assumption 5’, suppose that $G$ and $g$ satisfy
$G^{\prime}(g(\mu))=\mu\qquad\text{for all }\,\mu\geq 0.$
Then, a pair $(m,u)$ is a weak solution to Problem 1 in the sense of Definition 2.11 if and only if $u$ is an admissible minimizer of Problem 2 and $m=G^{\prime}(H(x,Du))$.
The detailed proof of this correspondence appears in Section 5.
Nonetheless, the variational formulation naturally gives rise to the contact-set condition on the boundary $\Gamma_{D}$, namely $(\psi-u)mD_{p}H(x,Du)\cdot\nu=0$. This condition, combined with the fact that the operator associated with the MFG (1.1)–(1.2) is monotone, helps us to prove the uniqueness result for the measure $m$, as well as a partial uniqueness result for the gradient of the value function, $u$.
Theorem 1.3.
Under Assumptions 1-5, any two pairs $(m,u),\,(\eta,v)\in L^{\beta+1}(\Omega)\times W^{1,\gamma}(\Omega)$ that solve Problem 1, in the sense of Definition 2.11, must satisfy the following.
  1. 1.
    For almost all $x\in\Gamma_{D}$, we have that
    $\begin{split}(\psi-u)\,\eta\,D_{p}H(x,Dv)\cdot\nu=(\psi-v)\,m\,D_{p}H(x,Du)\cdot\nu=0.\end{split}$
  2. 2.
    The equality $\eta(x)=m(x)$ holds for almost all $x\in\Omega$.
  3. 3.
    The equality $Du=Dv$ for almost all $E_{+}(m):=\{x\in\Omega\,:\,m(x)>0\}$.
The proofs of these uniqueness results are detailed in Section 6, while the uniqueness of minimizers for Problem 2 is addressed in Section 6.3. The paper concludes by discussing the Neumann trace in Appendix A.
2 Assumptions, Preliminary Results, and Weak Solutions
In this section, we lay out our notation, prescribe the assumptions used to prove our results, and give a suitable definition of weak solutions for Problem 1. These assumptions ensure the well-posedness of the variational formulation and enable us to define the boundary conditions in the MFG rigorously.
2.1 Notation
Given an exponent $\tau>1$, $\tau^{\prime}$ denotes its convex (Hölder) conjugate; that is,
$\dfrac{1}{\tau}+\dfrac{1}{\tau^{\prime}}=1.$
A differential operator written without a subscript (e.g., $Du$ or $\text{div}F$) refers to differentiation with respect to the variable $x$. Differentiation with respect to other variables will be indicated by a corresponding subscript. For example, $D_{p}H(x,p)$ denotes differentiation with respect to the second argument, $p$, even when $p$ is replaced by $Du$. For a quantity $f$, $f_{+}$ and $f_{-}$ are the positive and negative parts of $f$.
Throughout, we adopt the convention that any inequality involving a constant $C$ holds for any sufficiently large $C>0$ independent of the free variables; thus, replacing $C$ with a larger constant does not affect the validity of the inequality. If the constant depends on any free variable (e.g., $m$), this dependence is indicated with a suitable subscript (e.g., $C_{m}$) while maintaining the same interpretation. This convention minimizes unnecessary notation and allows for $C$ to vary from one line to another.
2.2 Assumptions
We begin by introducing the exponents we use to formulate our main assumptions and results. Assumption 1 concerns several parameters of the problem, while Assumptions 2 and 3 address the regularity of the boundary data $j$ and $\psi$. Assumptions 4 and 5 are typical assumptions in this setting on the growth of $H$, its derivatives, and $g$. Finally, Assumption 5’ concerns the function $G$ in the variational problem, Problem 2. The correspondence between Assumption 5 and 5’ is examined in Subsection 2.3.
Assumption 1.
Our assumptions are formulated in terms of two key exponents:
$\alpha>1\,\text{ and }\,\beta>0.$ (2.1)
To simplify the presentation, we introduce the auxiliary exponent:
$\gamma:=\dfrac{\beta+1}{\beta}\alpha.$ (2.2)
The data in the Neumann and Dirichlet boundary conditions satisfy the following assumptions.
Assumption 2.
The incoming flow $j\in L^{\gamma^{\prime}}(\Gamma_{N})$ is nonnegative, and does not vanish identically; that is, $j\geq 0$ and $j\not\equiv 0$ on $\Gamma_{N}$.
[…]
Assumption 3.
The exit cost is the trace of a function $\psi\in W^{1,\gamma}(\Omega)$ defined on the whole space.
Remark.
  1. $\diamond$
    The assumption that $\psi$ is defined on the whole domain $\Omega$ is useful for establishing a lower bound for the variational functional.
  2. $\diamond$
    By standard trace theorems, this assumption is equivalent to $\psi\in W^{1-\frac{1}{\gamma},\gamma}(\Omega)$, which imposes minimal regularity for the well-posedness of our problem.
The following assumption pertains to the Hamiltonian $H$, which governs the interior dynamics of the MFG (Problem 1) and also appears in the variational formulation (Problem 2).
Assumption 4.
The Hamiltonian $H:\mkern 1.0mu\overline{\mkern-1.0mu\Omega\mkern-1.0mu}\mkern 1.0mu\times\mathbb{R}^{d}\to\mathbb{R}$ satisfies the following conditions.
  1. i.
    The maps $(x,p)\mapsto H(x,p)$ and $(x,p)\mapsto D_{p}H(x,p)$ are measurable in $x$ and continuous with respect to $p$.
  2. ii.
    For almost every $x\in\Omega$, the map $p\mapsto H(x,p)$ is strictly convex.
  3. iii.
    There is a constant $C>1$ such that, for almost all $x\in\Omega$ and all $p\in\mathbb{R}^{d}$,
    $C^{-1}\,|p|^{\alpha}-C\leq H(x,p)\leq C|p|^{\alpha}+C.$
  4. iv.
    There is a constant $C>1$ such that, for almost all $x\in\Omega$ and all $p\in\mathbb{R}^{d}$,
    $|D_{p}H(x,p)|\leq C(|p|^{\alpha-1}+1).$
  5. v.
    There is a constant $C>1$ such that, for almost all $x\in\Omega$ and all $p\in\mathbb{R}^{d}$,
    $-H(x,p)+D_{p}H(x,p)\cdot p\geq C^{-1}\,|p|^{\alpha}-C.$
Remark.
  1. $\diamond$
    These conditions ensure that the integrands appearing in our treatment are well-defined and properly contained in suitable $L^{p}$-spaces.
  2. $\diamond$
    Assumption 4.v provides a link to the underlying optimal control problem, whose Lagrangian is retrieved by the Legendre transform
    $L(x,v)=\sup_{p}\big{(}-v\cdot p-H(x,p)\big{)}.$
    When the control $v$ is optimal, it has the form $v^{*}=-D_{p}H(x,p)$. Hence,
    $L(x,v^{*})=D_{p}H(x,p)\cdot p-H(x,p),$
    is coercivity according to our assumption.
  3. $\diamond$
    An example of a Hamiltonian that satisfies the assumptions is
    $H(x,p)=(|p|^{2}+1)^{\alpha/2}+1.$
The following assumption pertains to the coupling term $g$, which appears in the MFG system (Problem 1) as a density-dependent potential that deters agents from high-density regions.
Assumption 5.
The coupling term $g$ satisfies the following conditions.
  1. i.
    The map $g:[0,\infty)\to\mathbb{R}$ is continuous and strictly increasing.
  2. ii.
    There is constant $C>1$ such that
    $\frac{m^{\beta}}{C}-C\leq g(m)\leq Cm^{\beta}+C\quad\text{ for all}\enskip m\in[0,\infty).$
Remark.
  1. $\diamond$
    These conditions ensure that the integrands appearing in our treatment are well-defined and properly contained in suitable $L^{p}$–spaces.
  2. $\diamond$
    The strict monotonicity assumption makes $g$ invertible, allowing us to establish the variational formulation of the MFG problem.
  3. $\diamond$
    An example of a coupling term that satisfies the assumptions is
    $g(m)=(m+1)^{\beta}.$ (2.3)
For the treatment of the variational problem, it is convenient to express our assumptions explicitly in terms of $G$. The following assumptions are equivalent to Assumption 5 above due to the Definition 2.1, which we state after the assumption. This equivalence is established in Lemma 2.3.
Assumption 5’.
There is a constant $z_{0}\in\mathbb{R}$ such that $G$ satisfies the following conditions.
  1. i.
    The map $G$ is constant for all $z\leq z_{0}$; that is, $G(z)=G(z_{0})$ for all $z\leq z_{0}$.
  2. ii.
    The map $G:\mathbb{R}\to\mathbb{R}$ is continuously differentiable.
  3. iii.
    The map $G$ is strictly convex on $[z_{0},\infty)$.
  4. iv.
    The map $G$ is strictly increasing on $[z_{0},\infty)$.
  5. v.
    There is a constant $C>0$ such that, for all $z\in[z_{0},\infty)$,
    $C^{-1}\,z_{+}^{1/\beta}-C\leq G^{\prime}(z)\leq C\,z_{+}^{1/\beta}+C.$
Remark.
  1. $\diamond$
    Note that Assumption 5’.v implies that there is a constant $C>0$ such that, for all $z\in[z_{0},\infty)$,
    $C^{-1}\,z_{+}^{(\beta+1)/\beta}-C\leq G(z)\leq C\,z_{+}^{(\beta+1)/\beta}+C.$ (2.4)
  2. $\diamond$
    Convexity and lower-boundedness, which we obtain from Assumption 5’.i and (2.4), respectively, are standard assumptions to ensure the existence of a minimizer for the functional (1.9).
  3. $\diamond$
    An example of a map $G$ that satisfies the assumptions is
    $G(z)=\begin{cases}\frac{\beta}{\beta+1}z^{(\beta+1)/\beta}-z+c&z\geq 1\\ c-\frac{1}{\beta+1}&z<1,\end{cases}$ (2.5)
    for some arbitrary constant $c\in\mathbb{R}$.
2.3 Equivalence between Assumption 5 and Assumption 5’
The equivalence between Assumption 5 and Assumption 5’ is crucial to ensure the correspondence between Problem 1 and Problem 2 given in Theorem 1.2. This equivalence depends on the relation
$G^{\prime}(z)=g^{-1}(z),$ (2.6)
where $g^{-1}(z)$ is the extended inverse of $g$ defined in the following.
Definition 2.1.
Let $g:[0,\infty)\to\mathbb{R}$ satisfy Assumption 5. We say that $g^{-1}:\mathbb{R}\to[0,\infty)$ the extended inverse of $g$ if,
  1. i.
    $g(g^{-1}(z))=z$   for all $z>g(0)$, and
  2. ii.
    $g^{-1}(z)=0$   for all $z\leq g(0)$.
Remark 2.2.
  1. $\diamond$
    Note that the exponent governing the growth of $G$ is $(\beta+1)^{\prime}=(\beta+1)/\beta$.
  2. $\diamond$
    Note also that $z_{0}=g(0)$ in Assumption 5’.
Lemma 2.3.
Let $G$ be defined according to (2.6). Then, Assumption 5 and Assumption 5’ are equivalent.
Proof.
Step 1 (5 $\implies$ 5’): First, suppose that $g$ satisfies Assumption 5, and set $z_{0}=g(0)$.
  1. i.
    By Definition 2.1, we have that $G^{\prime}(z)=g^{-1}(z)=0$ for all $z\leq z_{0}$. Hence, $G$ is constant on $(-\infty,z_{0}]$. In particular,
    $G(z)=G(z_{0})\quad\text{ for all }\,z\leq z_{0}.$
  2. ii.
    Because $G^{\prime}=g^{-1}\in C^{0}(\mathbb{R})$, we have $G\in C^{1}(\mathbb{R})$.
  3. iii.
    Because $g$ is increasing, its inverse $G^{\prime}$ is also increasing on $[z_{0},\infty)$. Therefore, $G$ is convex.
  4. iv.
    Because $g\bigl{(}[0,\infty)\bigr{)}=[z_{0},\infty)$, for every $z>z_{0}$ there is an $m>0$ such that $z=g(m)$. Consequently,
    $G^{\prime}(z)=G^{\prime}\bigl{(}g(m)\bigr{)}=m>0.$
    Hence, $G$ is strictly increasing on $\bigl{[}z_{0},\infty\bigr{)}$.
  5. v.
    For an arbitrary $z>z_{0}$, substituting $m=G^{\prime}(z)$ into Assumption 5.ii gives us
    $\frac{G^{\prime}(z)^{\beta}}{C}-C\leq g(G^{\prime}(z))=z\leq C\,G^{\prime}(z)^{\beta}+C.$ (2.7)
    The left-hand side gives us that
    $G^{\prime}(z)\leq(Cz+C)^{1/\beta}\leq(Cz_{+}+C)^{1/\beta}\leq Cz_{+}^{1/\beta}+C.$
    Notice that $C\,G^{\prime}(z)^{\beta}+C\geq 0$. Thus, the right-hand side in (2.7) gives us that
    $z_{+}^{1/\beta}\leq\left(C\,G^{\prime}(z)^{\beta}+C\right)^{1/\beta}\leq C\,G^{\prime}(z)+C.$
    The two inequalities above imply that
    $\frac{z_{+}^{1/\beta}}{C}-C\leq G^{\prime}(z)\leq Cz_{+}^{1/\beta}+C.$ (2.8)
Step 2 (5’ $\implies$ 5): To establish the converse, suppose that $G$ satisfies Assumption 5’ and note that $g=(G^{\prime})^{-1}$ immediately gives us that $g(0)=z_{0}$.
  1. i.
    Because $G^{\prime}$ is continuous, coercive, and strictly increasing on $[g(0),\infty)$, its inverse $g$ is continuous and strictly increasing on $[0,\infty)$.
  2. ii.
    For an arbitrary $m\geq 0$, substituting $z=g(m)$ into Assumption 5’.v gives us
    $\frac{(g(m)_{+})^{1/\beta}}{C}-C\leq G^{\prime}(g(m))=m\leq C(g(m)_{+})^{1/\beta}+C.$
    With some manipulations similar to the above, we obtain
    $\frac{m^{\beta}}{C}-C\leq g(m)\leq Cm^{\beta}+C.$
This concludes the proof. ∎
2.3.1 An auxiliary bound for $G$.
The following lemma enables us to calculate the Gateaux derivative in Section 5.
Lemma 2.4.
Suppose that Assumptions 4 and 5’ hold, and let $p_{0},p_{1}\in\mathbb{R}^{d}$ be arbitrary. There is a constant $C>0$, independent of $p_{0}$ and $p_{1}$ such that
$\frac{1}{|\varepsilon|}\lvert G\left(H(x,p_{0}+\varepsilon p_{1})\right)-G\left(H(x,p_{0})\right)\rvert\leq C|p_{0}|^{\gamma}+C|p_{1}|^{\gamma}+C$
for all $\varepsilon\in(-1,1)$.
Proof.
Let $\varepsilon\in(-1,1)$ be arbitrary. By the mean value theorem, for every $x\in\Omega$, there is a $\theta_{x,\varepsilon}\in(0,1)$ such that
$\begin{split}\frac{1}{\varepsilon}\big{[}G(H(x,p_{0}+\varepsilon p_{1}))&-G(H(x,p_{0}))\big{]}=G^{\prime}(H(x,p_{0}+\varepsilon\theta_{x,\varepsilon}p_{1}))\,\,D_{p}H(x,p_{0}+\varepsilon\theta_{x,\varepsilon}p_{1})\cdot p_{1}.\end{split}$ (2.9)
Assumption 4.iv and Young’s inequality give us that
$\begin{split}|D_{p}H(x,p_{0}+\varepsilon\theta_{x,\varepsilon}p_{1})\cdot p_{1}|&\leq C|p_{0}|^{\alpha}+C|p_{1}|^{\alpha}.\end{split}$ (2.10)
Similarly, from the monotonicity of $G^{\prime}$, along with Assumption 5’.v and Assumption 4.iii, we deduce
$G^{\prime}(H(x,p_{0}+\varepsilon\theta_{x,\varepsilon}p_{1}))\leq C|p_{0}|^{\alpha/\beta}+C|p_{1}|^{\alpha/\beta}+C.$ (2.11)
By combining (2.9)–(2.11) and applying Young’s inequality, we obtain
$\frac{1}{|\varepsilon|}\lvert G\left(H(x,p_{0}+\varepsilon p_{1})\right)-G\left(H(x,p_{0})\right)\rvert\leq C|p_{0}|^{\alpha(\beta+1)/\beta}+C|p_{1}|^{\alpha(\beta+1)/\beta}+C.$
We conclude by recalling that $\gamma=\alpha(\beta+1)/\beta$. ∎
2.4 Formal estimates
Lemma 2.5.
Suppose Assumptions 2 and 3 hold. Let $u\in\mathcal{U}$ be an admissible function in Problem 2. Then, there is a constant $C>1$, independent of $u$, such that
$\int_{\Gamma_{N}}ju{\,\,\rm d}s\leq C\|Du\|_{L^{\gamma}(\Omega)}+C.$
Proof.
First, using Hölder’s inequality and that $(u-\psi)_{+}\equiv 0$ on $\Gamma_{D}$, we obtain
$\begin{split}\int_{\Gamma_{N}}j(u-\psi)_{+}&\leq\|j\|_{L^{\gamma^{\prime}}(\Gamma_{N})}\enskip\|(u-\psi)_{+}\|_{L^{\gamma}(\Gamma_{N})}\\ &=C\,\|(u-\psi)_{+}\|_{L^{\gamma}(\Gamma_{N}\cup\Gamma_{D})}.\end{split}$
The right-hand side can be further bounded by the trace theorem, giving us
$\int_{\Gamma_{N}}j(u-\psi)_{+}\leq C\|(u-\psi)_{+}\|_{L^{\gamma}(\Omega)}+C\|D(u-\psi)_{+}\|_{L^{\gamma}(\Omega)}.$
Moreover, because $(u-\psi)_{+}\equiv 0$ on $\Gamma_{D}$ and $|\Gamma_{D}|\not=0$ is predefined, we can apply the Poincaré inequality. This gives us
$\|(u-\psi)_{+}\|_{L^{\gamma}(\Omega)}\leq C\|D(u-\psi)_{+}\|_{L^{\gamma}(\Omega)}$
and, hence,
$\begin{split}\int_{\Gamma_{N}}j(u-\psi)_{+}&\leq C\|D(u-\psi)_{+}\|_{L^{\gamma}(\Omega)}\\ &\leq C\|Du\|_{L^{\gamma}(\Omega)}+C\|D\psi\|_{L^{\gamma}(\Omega)}.\end{split}$
Finally, because $j\geq 0$, we have
$\int_{\Gamma_{N}}j(u-\psi)\leq\int_{\Gamma_{N}}j(u-\psi)_{+}.$
Combining the inequalities above, we obtain
$\int_{\Gamma_{N}}j(u-\psi){\,\,\rm d}s\leq C\|Du\|_{L^{\gamma}(\Omega)}+C\|D\psi\|_{L^{\gamma}(\Omega)}.$
Our claim follows immediately. ∎
Now, we discuss some first-order apriori estimates for our problem. This class of estimates, well-known in MFGs, relies on the preceding lemma to control the boundary conditions. These formal estimates also motivate the functional spaces used in our proofs.
Proposition 2.6.
Under Assumptions 1-5, there is constant $C>0$, depending only on our data, such that, for any classical solution $(m,u)$ to Problem 1, we have
$\|m\|_{L^{\beta+1}(\Omega)}\leq C,\quad\|m|Du|^{\alpha}\|_{L^{1}(\Omega)}\leq C,\quad and\quad\|Du\|_{L^{\gamma}(\Omega)}\leq C.$
Proof.
Because $(m,u)$ solves the H-J equation in the system (1.1) pointwise, almost everywhere, we have that
$m=G^{\prime}(H(x,Du)).$
Subsequently, we also have
$C^{-1}|Du|^{\alpha/\beta}-C\leq m\leq C|Du|^{\alpha/\beta}+C.$
This inequality provides us with three auxiliary bounds:
$|Du|^{\gamma}\leq(Cm+C)^{\beta+1}\leq Cm^{\beta+1}+C,$ (2.12)
$m^{\beta+1}\leq(C|Du|^{\alpha/\beta}+C)^{\beta+1}\leq C|Du|^{\gamma}+C,$ (2.13)
and
$C^{-1}m^{\beta+1}-C\leq C^{-1}m(m-C)_{+}^{\beta}\leq m|Du|^{\alpha},$ (2.14)
where the constants appearing above are independent of $m$ and $u$. Subsequently,
$C^{-1}\|m\|_{L^{\beta+1}(\Omega)}^{\beta+1}-C\leq\|Du\|_{L^{\gamma}(\Omega)}^{\gamma}\leq C\|m\|_{L^{\beta+1}(\Omega)}^{\beta+1}+C.$
Now, testing the transport equation in (1.1) by $(u-\psi)$ gives us
$\begin{split}LHS=\int_{\Omega}m\,D_{p}H(x,Du)\cdot D(u-\psi){\,\,\rm d}x&=\int_{\Gamma}(u-\psi)\,m\,D_{p}H(x,Du)\cdot\nu{\,\,\rm d}s\\ &=\int_{\Gamma_{N}}j(u-\psi){\,\,\rm d}s.\end{split}$ (2.15)
The left-hand side (LHS) is bounded below due to Assumption 4 and Young’s inequality (with $\varepsilon$). More precisely, we have
$\begin{split}LHS&\geq\int_{\Omega}m\left(D_{p}H(x,Du)\cdot Du-|D_{p}H(x,Du)||D\psi|\right){\,\,\rm d}x\\ &\geq\int_{\Omega}m(C^{-1}|Du|^{\alpha}-\varepsilon|Du|^{\alpha}-C_{\varepsilon}|D\psi|^{\alpha}-C_{\varepsilon}){\,\,\rm d}x.\end{split}$
Selecting $\varepsilon$ appropriately gives us
$\begin{split}LHS&\geq C^{-1}\int_{\Omega}m|Du|^{\alpha}{\,\,\rm d}x-C\int_{\Omega}m(|D\psi|^{\alpha}+1){\,\,\rm d}x\\ \end{split}$
Using inequality (2.14) and Young’s inequality with $\varepsilon$, we obtain
$\begin{split}LHS&\geq C^{-1}\|m\|_{L^{\beta+1}(\Omega)}^{\beta+1}-\varepsilon\|m\|_{L^{\beta+1}(\Omega)}^{\beta+1}-C_{\varepsilon}\||D\psi|^{\alpha}+1\|_{L^{(\beta+1)^{\prime}}(\Omega)}^{(\beta+1)^{\prime}}\\ \end{split}$
Selecting $\varepsilon$ appropriately gives us
$\begin{split}LHS&\geq C^{-1}\|m\|_{L^{\beta+1}}^{\beta+1}-C\\ \end{split}$
Subsequently, (2.15) becomes
$\begin{split}\int_{\Gamma_{N}}j(u-\psi){\,\,\rm d}s&\geq C^{-1}\|m\|_{L^{\beta+1}}^{\beta+1}-C\\ \end{split}$
On the other hand, by Lemma 2.5 and Young’s inequality, we have
$\begin{split}C^{-1}\|m\|_{L^{\beta+1}}^{\beta+1}-C&\leq C\|Du\|_{L^{\gamma}(\Omega)}+C\\ &\leq\varepsilon\|Du\|_{L^{\gamma}(\Omega)}^{\gamma}+C_{\varepsilon}\leq\varepsilon C\|m\|_{L^{\beta+1}}^{\beta+1}+C\varepsilon\end{split}$
Selecting $\varepsilon$ appropriately, we obtain
$\begin{split}\|m\|_{L^{\beta+1}}^{\beta+1}\leq C.\end{split}$
Note that $C$ is independent of $m$ and $u$. The other two estimates follow immediately using (2.12) and (2.14). ∎
2.5 The normal trace theorem
Here, we present the normal trace theorem that ensures the well-definedness of the Neumann boundary condition in a suitable dual space. We briefly overview this theorem, deferring detailed proofs and discussion to Appendix A. To this end, we first define a function space tailored to our setting.
Definition 2.7.
The divergence Sobolev space $W^{\gamma^{\prime}}(\operatorname*{div};\Omega)$ is the space consisting of all vector fields
$\mathbf{w}\in L^{\gamma^{\prime}}(\Omega;\mathbb{R}^{d})\quad\text{such that}\quad\operatorname*{div}(\mathbf{w})\in L^{\gamma^{\prime}}(\Omega),$
where $\operatorname*{div}(\mathbf{w})$ is interpreted in the weak sense. (See Definition A.1 in the appendix). This space is, naturally, equipped with the norm
$\|\mathbf{w}\|_{W^{{\gamma^{\prime}}}(\operatorname*{div};\Omega)}:=\Big{(}\|\mathbf{w}\|_{L^{\gamma^{\prime}}(\Omega;\mathbb{R}^{d})}^{\gamma^{\prime}}+\|\operatorname*{div}(\mathbf{w})\|_{L^{\gamma^{\prime}}(\Omega)}^{\gamma^{\prime}}\Big{)}^{1/{\gamma^{\prime}}}.$
Now, we are ready to state the normal trace theorem.
Theorem 2.8.
Let $\Omega$ be an open, bounded, $C^{1}$–domain. Then, there exists a bounded linear operator
$\operatorname{\mathfrak{tr}}_{1}:W^{\gamma^{\prime}}(\operatorname*{div};\Omega)\to W^{-1+\frac{1}{\gamma},\,{\gamma^{\prime}}}(\Gamma),$
satisfying the following.
  1. (i)
    For all smooth vector fields $\mathbf{F}\in C^{\infty}(\mkern 1.0mu\overline{\mkern-1.0mu\Omega\mkern-1.0mu}\mkern 1.0mu;\mathbb{R}^{d})$,
    $\operatorname{\mathfrak{tr}}_{1}(\mathbf{F})=\mathbf{F}|_{\Gamma}\cdot\nu.$
  2. (ii)
    There is a constant $C>0$ such that, for all $\mathbf{F}\in W^{\gamma^{\prime}}(\operatorname*{div};\Omega)$,
    $\|\operatorname{\mathfrak{tr}}_{1}(\mathbf{F})\|_{W^{-1+\frac{1}{\gamma},{\gamma^{\prime}}}(\Gamma)}\leq C\|\mathbf{F}\|_{W^{\gamma^{\prime}}(\operatorname*{div};\Omega)}.$
Remark 2.9.
  1. i.
    The space $W^{-1+\frac{1}{\gamma},\,{\gamma^{\prime}}}(\Gamma)$ is the dual of $W^{1-\frac{1}{\gamma},\,{\gamma}}(\Gamma)$, which is the trace space of $W^{1,{\gamma}}(\Gamma)$ [DiB16].
  2. ii.
    Henceforth, we write $F\cdot\nu=\operatorname{\mathfrak{tr}}_{1}(F)$, where $\nu$ is the normal to the boundary $\Gamma$.
Proof.
We defer the proof to our discussion in the Appendix A. ∎
Remark 2.10.
The estimates from the previous section combined with Assumption 4 imply that $mD_{p}H(x,Du)\in L^{\gamma^{\prime}}(\Omega;\mathbb{R}^{d})$. Since $\operatorname*{div}(mD_{p}H(x,Du))=0$, we have $mD_{p}H(x,Du)\in W^{\gamma^{\prime}}(\operatorname*{div};\Omega)$. Thus, the previous theorem ensures that the normal trace $mD_{p}H(x,Du)\cdot\nu$ is well defined in $\Gamma_{N}$.
2.6 Definition of weak solutions
We define a weak solution by extending the definition employed in [CG15] to include additional constraints that rigorously account for boundary data. A key aspect of this definition is that we enforce boundary conditions in the sense of traces. More precisely, we define a weak solution as follows.
Definition 2.11.
A weak solution to Problem 1 is a pair of functions $(m,u)\in L^{\beta+1}(\Omega)\times W^{1,\gamma}(\Omega)$ that satisfies the following conditions.
  1. (C0)
    The density $m$ is nonnegative.
  2. (C1)
    The Hamilton-Jacobi equation in (1.1) holds in the following sense:
    $\begin{split}H(x,Du(x))=g(m(x)),\quad&\text{ a.e. in }\{x\in\Omega\colon m(x)>0\},\enskip\text{ and}\\ H(x,Du(x))\leq g(0),\quad&\text{ a.e. in }\{x\in\Omega\colon m(x)=0\}.\end{split}$
  3. (C2)
    The Dirichlet boundary condition in (1.2) is satisfied in the trace sense.
  4. (C3)
    The transport equation in (1.1) holds in the following sense.
    $\int_{\Omega}mD_{p}H(x,Du)\cdot D\phi{\,\,\rm d}x-\int_{\Gamma}\phi\,\,mD_{p}H(x,Du)\cdot\nu{\,\,\rm d}s=0$
    for all $\phi\in C_{c}^{\infty}(\mathbb{R}^{d})$, which are test functions defined on all of $\mathbb{R}^{d}$.
  5. (C4)
    The Neumann boundary conditions in (1.2) are satisfied in the Neumann trace sense.
3 Examples
This section presents explicit closed-form examples of MFGs that motivate our model and illustrate characteristic behaviors of solutions to (3.5)–(3.6). In particular, we demonstrate that imposing a strict Dirichlet boundary condition in first-order problems may lead to artificial inflows at the boundary, while phenomena such as the formation of empty regions (where $m\equiv 0$) and potential non-uniqueness of the value function $u$ can arise, issues of particular importance in scenarios with congestion or overcrowding. Moreover, these examples serve two key purposes: they provide benchmarks for verifying numerical methods, and they reveal intrinsic connections among the model’s components, thereby facilitating the development of new theoretical results (see, e.g., [Bar12]).
3.1 Limitations of Strict Dirichlet Conditions in Exit Problems
In exit problems in optimal control, the system’s behavior within the domain is modeled by a Hamilton–Jacobi equation, while the exit cost is prescribed via a Dirichlet boundary condition. In the context of viscosity solutions, the boundary value may not be achieved, and the solution becomes discontinuous at the boundary. A simple example is the Hamilton-Jacobi equation
$\frac{1}{2}u_{x}^{2}=0\quad\text{in }(0,1),$
subjected to the Dirichlet boundary conditions $u(0)=0$ and $u(1)=1$. The unique viscosity solution for this problem is $u=0$ on $(0,1)$, which is discontinuous at $x=1$. However, in the context of MFGs, strictly imposing the Dirichlet condition (i.e., specifying an equality rather than an inequality) can lead to unintended outcomes as discontinuities at the boundary may not arise and instead lead to unintended inflows of agents.
To illustrate this behavior, consider the MFG
$\begin{cases}\frac{1}{2}u_{x}^{2}=m&\text{in }(0,1),\\[2.84526pt] -(m\,u_{x})_{x}=1&\text{in }(0,1),\end{cases}$ (3.1)
subject to the Dirichlet boundary conditions
$u(0)=0\quad\text{and}\quad u(1)=a.$
Note that the source term in the second equation in (3.1) obviates the need for an inflow of agents, thus emphasizing the role of the Dirichlet boundary condition.
The system is explicitly solvable. Integrating the second equation in (3.1) yields that the flux of agents is
$j(x):=-m\,u_{x}=x+c,$ (3.2)
for some constant $c\in\mathbb{R}$. Substituting this expression into the Hamilton–Jacobi equation, $\frac{1}{2}u_{x}^{2}=m$, leads to
$m=\frac{1}{2^{1/3}}(x+c)^{2/3}.$
Next, substituting back into (3.2) and applying the boundary condition $u(0)=0$ yields
$u(x)=\frac{3}{2^{5/3}}\Big{[}c^{4/3}-(x+c)^{4/3}\Big{]}.$
Finally, imposing the boundary condition at $x=1$ determines $c$ via
$\frac{3}{2^{5/3}}\Big{[}c^{4/3}-(1+c)^{4/3}\Big{]}=a.$ (3.3)
Applying the mean value theorem to the difference $(1+c)^{4/3}-c^{4/3}$ yields
$a=-2^{1/3}\Big{(}c+\theta_{c}\Big{)}^{1/3},$ (3.4)
with some $\theta_{c}\in(0,1)$. Consequently, a large positive $a$ forces $c$ to be significantly negative, while a large negative $a$ results in a correspondingly large positive $c$.
Now, we examine how variations in $a$ influence the flow direction at the boundaries. The flux at the boundaries is given by
$j(0)=c\quad\text{and}\quad j(1)=1+c.$
Recall that the interpretation of these fluxes depends on the outward normal: at $x=0$, a negative value $j(0)$ indicates outflow (exit), whereas at $x=1$, a positive value $j(1)$ indicates outflow. Thus, if $a$ is sufficiently large and positive (so that $c\ll-1$), then $j(0)$ is negative—implying an exit at $x=0$, while $j(1)$ corresponds to an inflow (entry). Conversely, if $a$ is very negative, then $c$ becomes positive, so that $j(0)$ indicates an entry at $x=0$ and $j(1)$ an exit at $x=1$. A third scenario arises when $-1<c<0$; in this case, the fluxes are such that both boundaries effectively serve as exit points. The reason for this behavior is the following. Since the running cost scales with the density of players, imposing an excessively high Dirichlet boundary condition may force the system to achieve an unrealistically high density near the boundary. Consequently, the system may induce an artificial inflow of players to satisfy the boundary requirement, as seen in our example, where a large positive $a$ resulted in $x=1$ functioning as an entry point.
In summary, these observations confirm that the Dirichlet boundary conditions do not predetermine the entry-exit regions of the domain boundary; one must solve the problem to determine the actual behavior. This motivates the adoption of a relaxed Dirichlet condition to ensure that exit and entry regions are correctly and a priori identified.
3.2 One-Dimensional Mean Field Games with Mixed Boundary Conditions
In contrast to the strict Dirichlet boundary condition from Subsection 3.1, we now consider a simple one-dimensional MFG with a relaxed Dirichlet condition at one end of the interval and a prescribed Neumann inflow at the other. This example admits an explicit solution that offers insights into the interplay between inflow constraints, exit costs, and the resulting player density of MFG systems of the form (1.1)–(1.2).
Let $\Omega=(0,1)$, and consider the MFG system
$\begin{cases}\frac{1}{2}u_{x}^{2}+V(x)=m,\\ -(m\,u_{x})_{x}=0,\end{cases}$ (3.5)
subject to the boundary conditions
$\begin{cases}-m(0)\,u_{x}(0)=j_{0},\\[2.5pt] u(1)\leq 0,\\[2.5pt] u(1)\,m(1)\,u_{x}(1)=0,\end{cases}$ (3.6)
where the nonnegative constant $j_{0}$ corresponds to the prescribed inflow of players at $x=0$. The key distinction of this setup is that we impose a relaxed Dirichlet condition at $x=1$, which specifies the players’ exit cost but allows for selective departure. This strategic combination of boundary conditions is crucial for the precise modeling of the flow of players within the domain and for ascertaining the influence of exit costs on the players’ flow and distribution. Moreover, the term $V$ is called the potential, and it represents the spatial preferences of the players in MFG models (i.e., distribution of attraction sites in the domain of the game).
We divide our analysis of this example into two cases: $j_{0}=0$ (no inflow) and $j_{0}>0$ (positive inflow).
Case 1 ($j_{0}=0$):
In the absence of inflow, the transport equation in (3.5) and the left boundary condition in (3.6) imply that
$m(x)u_{x}(x)=m(0)u_{x}(0)=0\quad\text{for all }x\in(0,1);$
hence, either $u_{x}(x)=0$ or $m(x)=0$. Taking Definition 2.11 into consideration, we substitute these alternatives into the HJ equation in (3.5). Thus, we obtain that either
$\begin{array}[]{ll}m(x)>0\enskip\text{ and }\enskip u_{x}(x)=0&\implies m(x)=V(x),\end{array}$
or
$\begin{array}[]{ll}m(x)=0&\implies\frac{1}{2}u_{x}(x)^{2}\leq-V(x).\end{array}$
The first alternative is only possible if $m(x)=V(x)=V_{+}(x)$, while the second alternative can only hold if $V\leq 0$, implying also that $m(x)=V_{+}(x)=0$. A particular solution can be obtained by considering the equality case in the Hamilton-Jacobi equation, leading to the solution
$\begin{cases}m(x)=V_{+}(x),\\[5.0pt] u(x)=u(1)\pm\displaystyle\sqrt{2}\int_{x}^{1}\sqrt{V_{-}(z)}{\,\,\rm d}z.\end{cases}$
Thus, if $V$ changes sign, the MFG (3.5)–(3.6) admits at least two distinct solutions for $u$.
This example illustrates that the density $m$ may vanish in certain regions. Owing to the invariance of $u$ under the addition of a constant, we may assume without loss of generality that $u(1)=0$. In the zero-current case, either the velocity $u_{x}$ is zero (so that $m\geq 0$) or the velocity is nonzero, in which case $m$ must vanish.
Case 2 ($j_{0}>0$):
From the Fokker-Plank equation and the boundary condition, we must have
$-m(x)u_{x}(x)=m(0)u_{x}(0)=j_{0}.$
Because $j_{0}\neq 0$, it follows that $m(x)>0$ for all $x$; hence,
$u_{x}(x)=-\frac{j_{0}}{m(x)}.$ (3.7)
Substituting this into the first equation of (3.5) and multiplying by $m(x)^{2}$ yields
$m(x)^{3}-V(x)\,m(x)^{2}-\frac{j_{0}^{2}}{2}=0.$ (3.8)
This cubic equation admits three real roots, of which exactly one is positive. Thus, we obtain
$\begin{split}m(x)=\begin{cases}\sqrt[3]{a(x)+b(x)}+\sqrt[3]{a(x)-b(x)}+\dfrac{V(x)}{3},&V>\gamma,\\ \dfrac{j_{0}^{2/3}}{2},&V(x)=\gamma,\\ v^{k_{0}}\sqrt[3]{a(x)+b(x)}+\dfrac{1}{v^{k_{0}}\sqrt[3]{a(x)+b(x)}}\left(\dfrac{V}{3}\right)^{2}+\dfrac{V}{3},&V(x)<\gamma,\end{cases}\end{split}$
where $v$ is a complex cube root of $1$, $k_{0}=0,1,2$, and
$\begin{split}\gamma=-\dfrac{3}{2}j_{0}^{2/3},\quad a(x)=\left(\frac{V}{3}\right)^{3}+\left(\frac{j_{0}}{2}\right)^{2},\quad b(x)=\frac{j_{0}}{2}\sqrt{2\left(\frac{V}{3}\right)^{3}+\left(\frac{j_{0}}{2}\right)^{2}}.\end{split}$
Although we have a complex expression in the case of $V(x)<\gamma$, all three roots of $m$, for $k_{0}=0,1,2$, are real due to the coefficients of the cubic equation. Moreover, two of them are negative, and only one is positive, which we choose.
From (3.7), it follows that $u_{x}(x)<0$. Consequently, the first equation in (3.5) implies
$u_{x}(x)=-\sqrt{2\bigl{(}m(x)-V(x)\bigr{)}},$
where the negative square root is chosen for consistency. Imposing the compatibility condition $u(1)=0$ from (3.6) yields
$u(x)=\int_{x}^{1}\sqrt{2\bigl{(}m(x)-V(x)\bigr{)}}{\,\,\rm d}x.$
This example illustrates that, in one dimension, a positive current guarantees that the density $m$ remains strictly positive—regardless of the sign and amplitude of the potential $V$. This behavior contrasts with the two-dimensional case, where $m$ may vanish in certain regions, as we discuss in Subsection 3.3.
3.2.1 Confirmation of the variational formulation
It is worth noting that, in this example, the variational formulation that corresponds to the MFG (3.5)–(3.6) is given by
$\inf_{u}\int_{0}^{1}\frac{1}{2}\left(\Bigl{(}\frac{1}{2}u_{x}^{2}+V(x)\Bigr{)}_{+}\right)^{2}{\,\,\rm d}x-j_{0}u(1),$ (3.9)
subject to the constraint $u(1)\leq 0$. Here, the coupling term $g(m)=m$ and
$G(z)=\frac{1}{2}(z_{+})^{2}.$
This variational problem can be solved numerically via a simple finite difference method in conjunction with Mathematica’s built-in function FindMinimum. As a result, one can obtain an experimental confirmation of the correspondence result in Theorem 1.2.
The case $j_{0}=0$:
Figure 1 compares the numerical results with the analytical solution.
Figure 1: Comparison of the analytical (red dashed) and numerical (green) solutions for $m$ and two distinct $u_{1}$ and $u_{2}$, with $j_{0}=0$, and $V(x)=0.5\sin\bigl{(}3\pi(x+0.25)\bigr{)}$.
The case $j_{0}>0$:
Figure 2 compares the numerical results and the analytical solution.
Figure 2: Comparison of the analytical (red dashed) and numerical (green) solutions for $u$ and $m$, with $j_{0}=0.1$ and $V(x)=\gamma+0.5\sin\bigl{(}3\pi(x+0.25)\bigr{)}$.
3.3 A two-dimensional example
Let $\Omega=(0,1)\times(0,1)$ be the unit square whose boundary is partitioned as $\Gamma=\Gamma_{N}\cup\Gamma_{D}$, where
$\begin{split}\Gamma_{N}&=\{0\}\times[0,1]\cup[0,1]\times\{1\},\quad\text{and}\\ \Gamma_{D}&=\{1\}\times[0,1]\cup[0,1]\times\{0\}.\end{split}$
Set
$V(x,y):=3e^{-\pi x}\cos(\pi y)-\frac{1}{2}\pi^{2}e^{-2\pi x},$
and consider the following MFG problem on $\Omega$:
The MFG System:
$\begin{cases}\frac{1}{2}|Du(x,y)|^{2}+V(x,y)=m,&\text{for }(x,y)\in\Omega,\\[2.84526pt] -\operatorname*{div}(mDu)=0,&\text{for }(x,y)\in\Omega.\end{cases}$
Neumann Boundary Conditions:
$\begin{cases}-m\,\partial_{x}u(0,y)=\dfrac{3\pi}{2}\,[\sin(2\pi y)]_{+},&\text{for }y\in[0,1],\\[3.50006pt] m\,\partial_{y}u(x,1)=0,&\text{for }x\in[0,1].\end{cases}$
Relaxed Dirichlet Boundary Conditions:
$\begin{cases}u(1,y)\leq e^{-\pi}\sin(\pi y),&\text{for }y\in[0,1],\\[3.50006pt] u(x,0)\leq 0,&\text{for }x\in[0,1].\end{cases}$
No-Entry Conditions:
$\begin{cases}m\,\partial_{x}u(1,y)\leq 0,&\text{if }u(1,y)=e^{-\pi}\sin(\pi y)\text{ for }y\in[0,1],\\[3.50006pt] -m\,\partial_{y}u(x,0)\leq 0,&\text{if }u(x,0)=0\text{ for }x\in[0,1].\end{cases}$
Contact-Set Conditions:
$\begin{cases}m\bigl{(}u(1,y)-e^{-\pi}\sin(\pi y)\bigr{)}\,\partial_{x}u(1,y)=0,&\text{for }y\in[0,1],\\[3.50006pt] -m\,u(x,0)\,\partial_{y}u(x,0)=0,&\text{for }x\in[0,1].\end{cases}$
A direct computation shows that the pair
$\begin{cases}u(x,y)=e^{-\pi x}\sin(\pi y),\\[2.84526pt] m(x,y)=3e^{-\pi x}\,[\cos(\pi y)]_{+},\end{cases}$
solves the above MFG in the sense of Definition 2.11.
(a) Graph of $m$.
(b) Graph of the flow, $J$, where the upper half shows the absence of flow in empty regions.
Figure 3: Illustration of MFG Behavior: Density $m$ and Flux $J$ Highlighting Empty Regions and Flow Patterns.
This example illustrates several intrinsic features of the MFG. We highlight two phenomena relevant to our discussion:
  1. 1.
    In the above example, the function $\psi$ on the Dirichlet boundary is given by
    $\psi(x,y)=\begin{cases}e^{-\pi}\sin(\pi y)&\text{for }x=1,\;y\in[0,1],\\[3.50006pt] 0&\text{for }x\in[0,1],\;y=0;\end{cases}$
    Accordingly, $u=\psi$ on $\Gamma_{D}$. However, the constraints imply
    $\begin{cases}m\,\partial_{x}u(1,y)=-\dfrac{3\pi e^{-2\pi}}{2}\,[\sin(2\pi y)]_{+}\leq 0,&\text{for }y\in[0,1],\\[3.50006pt] -m\,\partial_{y}u(x,0)=-3\pi e^{-2\pi x}\,[\cos(\pi y)]_{+}\leq 0,&\text{for }x\in[0,1].\end{cases}$
    In particular, since $m\,\partial_{x}u(1,y)=0$ for $y\in[\tfrac{1}{2},1]$, the condition $u=\psi$ on $\Gamma_{D}$ does not necessarily force players to exit the domain in that region.
  2. 2.
    Although in this example $u=\psi$ on $\Gamma_{D}$, one may modify the boundary datum by setting
    $\widetilde{\psi}(x,y)=\begin{cases}e^{-\pi}\sin(\pi y)+\bigl{(}y-\tfrac{1}{2}\bigr{)}_{+}^{2}&\text{for }x=1,\;y\in[0,1],\\[3.50006pt] 0&\text{for }x\in[0,1],\;y=0,\end{cases}$
    thus, yielding a modified problem that still admits the same solution. This modification illustrates that a solution $u$ may differ from $\psi$ on $\Gamma_{D}$, emphasizing the necessity for the relaxation of the Dirichlet boundary condition.
3.4 More two-dimensional examples with a complex variable construction
We introduce a novel method for generating explicit solutions to MFGs using properties of complex-valued functions in two dimensions. This approach enables the construction of solutions that exhibit regions devoid of agents, which is the first time such a method has been proposed in the MFG literature. The method takes advantage of the fact that any harmonic function in a connected domain admits a harmonic conjugate, with the potential $V$ chosen retrospectively to satisfy the MFG system.
More precisely, we consider MFGs of the form
$\begin{split}\begin{cases}\frac{1}{2}|Du|^{2}+V=m^{1/q}&\text{ for }(x,y)\in\Omega,\\ -\operatorname*{div}(mDu)=-Dm\cdot Du-m\Delta u=0&\text{ for }(x,y)\in\Omega.\end{cases}\end{split}$ (3.10)
Consider a differentiable, complex-valued mapping
$z=x+iy\mapsto f(z)=u(x,y)+i\,m(x,y),$
where $u$ and $m$ are real-valued functions. By the Cauchy-Riemann equations, we have
$\begin{split}\begin{cases}\partial_{x}u&=\partial_{y}m\\ \partial_{y}u&=-\partial_{x}m.\end{cases}\end{split}$
Thus,
$\begin{split}Du\cdot Dm=\partial_{x}u\,\partial_{x}m+\partial_{y}u\,\partial_{y}m=0.\end{split}$
Because $f$ is differentiable, $u$ is harmonic; that is, $u$ solves $-\Delta u=0$. Then, $m$ and $u$ solve the continuity equation;
$\begin{split}-\operatorname*{div}(mDu)=-Dm\cdot Du-m\Delta u=0.\end{split}$ (3.11)
We then choose the potential $V(x,y)$ so that the first equation in (3.10) holds:
$\begin{split}V(x,y)=m(x,y)^{1/q}-\frac{1}{2}|Du(x,y)|^{2}\end{split}$
Generally, the function $m$ obtained from this construction may not be nonnegative everywhere, which is problematic for a density function. To remedy this issue, we introduce a transitional adjustment by replacing $m$ with its positive part, denoted by $\widetilde{m}=(m)_{+}$, and define,
$\begin{split}V(x,y)=\widetilde{m}(x,y)^{1/q}-\frac{1}{2}|Du(x,y)|^{2}.\end{split}$
Then, we have
$\begin{split}\begin{cases}\frac{1}{2}|Du(x,y)|^{2}+V(x,y)=\widetilde{m}^{1/q}&\text{ for }(x,y)\in\Omega,\\ -\operatorname*{div}(\widetilde{m}Du)=0&\text{ for }(x,y)\in\Omega.\end{cases}\end{split}$
The first equation holds in a weak-strong sense (by Definition 2.11, below), while the second equation holds in the sense of distributions.
4 The variational problem
In this section, we analyze the variational formulation in Problem 2 in more depth and prove the existence of minimizers. We additionally derive a few regularity estimates that arise directly from this formulation. We use the direct method in the calculus of variations, with the core difference lying in how we handle the nonstandard boundary conditions.
4.1 On the existence of a minimizer
In this subsection, we establish the existence of minimizes for Problem 2. As customary in the study of variational problems, we begin by establishing the following lower bound.
Proposition 4.1.
Under Assumptions 1-4 and Assumption 5’, let $u\in\mathcal{U}$ be an admissible function in Problem 2. Then, there is a constant $C>1$, independent of $u$, such that
$\mathcal{I}[u]\geq C^{-1}\|Du\|_{L^{\gamma}(\Omega)}^{\gamma}-C.$
Proof.
Let $u\in\mathcal{U}$ be arbitrary. Then, from Assumption 5’ and Assumption 4, we obtain
$\begin{split}G(H(x,Du))&\geq C^{-1}\,H(x,Du)_{+}^{(\beta+1)/\beta}-C\\ &\geq C^{-1}\,(|Du|^{\alpha}-C)_{+}^{(\beta+1)/\beta}-C\\ &\geq C^{-1}|Du|^{\gamma}-C,\end{split}$ (4.1)
where used that $\gamma=\alpha(\beta+1)/\beta$. This implies that,
$\begin{split}\mathcal{I}[u]&=\int_{\Omega}G(H(x,Du)){\,\,\rm d}x-\int_{\Gamma_{N}}ju{\,\,\rm d}s\\ &\geq C^{-1}\|Du\|_{L^{\gamma}(\Omega)}^{\gamma}-C-\int_{\Gamma_{N}}ju{\,\,\rm d}s.\end{split}$
Using Lemma 2.5, this gives us
$\begin{split}\mathcal{I}[u]&\geq C^{-1}\|Du\|_{L^{\gamma}(\Omega)}^{\gamma}-C\,\|Du\|_{L^{\gamma}(\Omega)}-C.\end{split}$
Since $\gamma>1$, this gives us that
$\mathcal{I}[u]\geq C^{-1}\|Du\|_{L^{\gamma}(\Omega)}^{\gamma}-C$
for some $C>0$, independent of $u$. ∎
Remark 4.2.
The previous proposition establishes that $\inf_{u\in\mathcal{U}}\,\mathcal{I}[u]$ is bounded below. It is also bounded above by
$\inf_{u\in\mathcal{U}}\,\mathcal{I}[u]\leq\mathcal{I}[\psi]=C.$
Next, we apply the direct method in the calculus of variations and prove that any minimizing sequence converges to an admissible minimizer, hence establishing the existence of a minimizer for the variational problem.
Proof of Theorem 1.1.
Let $\{u_{n}\}\in\mathcal{U}$ be a minimizing sequence. By Proposition 4.1, we have
$C^{-1}\|Du_{n}\|_{L^{\gamma}(\Omega)}^{\gamma}-C\leq\mathcal{I}[u_{n}]\leq C.$
Consequently,
$\|Du_{n}\|_{L^{\gamma}(\Omega)}\leq C$ (4.2)
for some $C>1$, independent of $n$. Moreover, due to Lemma 2.5 and Remark 4.2, we have that
$C^{-1}\|Du_{n}\|_{L^{\gamma}(\Omega)}-C\leq\int_{\Gamma_{N}}ju_{n}{\,\,\rm d}s\leq C\|Du_{n}\|_{L^{\gamma}(\Omega)}+C.$ (4.3)
And the extended Poincaré inequality implies that
$\|u_{n}\|_{L^{\gamma}(\Omega)}\leq C\|Du_{n}\|_{L^{\gamma}(\Omega)}+C\,\Big{|}\int_{\Gamma_{N}}ju_{n}{\,\,\rm d}s\Big{|}.$ (4.4)
Combining (4.2)–(4.4), we obtain that
$\|u_{n}\|_{L^{\gamma}(\Omega)}\leq C$ (4.5)
for some $C>1$, independent of $n$.
The estimates (4.2) and (4.5) give us that $\|u_{n}\|_{W^{1,\gamma}(\Omega)}\leq C$. Because $W^{1,\gamma}(\Omega)$ is a reflexive Banach space, the bounded sequence $\{u_{n}\}\subset W^{1,\gamma}(\Omega)$ is weakly precompact; that is, there exists a subsequence that converges weakly to a function $u\in W^{1,\gamma}(\Omega)$. We now show that
$u\in\mathcal{U}\quad\text{ and }\quad\liminf_{n\to\infty}\mathcal{I}[u_{n}]\geq\mathcal{I}[u].$
To this end, note that
$\mathcal{I}[u_{n}]-\mathcal{I}[u]=$ $\int_{\Omega}\left[G\left(H(x,Du_{n})\right)-G\left(H(x,Du)\right)\right]{\,\,\rm d}x-\int_{\Gamma_{N}}j\,(u_{n}-u){\,\,\rm d}s$
$\geq\int_{\Omega}G^{\prime}(H(x,Du))D_{p}H(x,Du)(Du_{n}-Du){\,\,\rm d}x-\int_{\Gamma_{N}}j(u_{n}-u){\,\,\rm d}s.$
Because $j\in L^{\gamma^{\prime}}(\Gamma_{N})$ and $G^{\prime}(H(\cdot,Du))D_{p}H(\cdot,Du)\in L^{\gamma^{\prime}}(\Omega)$, we have
[…]
and
$\begin{split}\int_{\Omega}G^{\prime}(H(x,Du))D_{p}H(x,Du)(Du_{n}-Du){\,\,\rm d}x\to 0.\end{split}$
It remains to show that $u\in\mathcal{U}$. Since $u_{n}\rightharpoonup u$ in $W^{1,\gamma}(\Omega)$, we have
[…]
for every $f\in L^{\gamma^{\prime}}(\Gamma_{D})$. Taking $f>0$, we conclude
[…]
Therefore, $u\leq\psi$ on $\Gamma_{D}$. ∎
4.2 First-order estimates
In this subsection, we establish several estimates that provide better regularity for solutions of the MFG. These estimates guarantee the applicability of some of our results (e.g., the Neumann trace theorem in the appendix).
Lemma 4.3.
Under Assumptions 1-4 and Assumption 5’, let $u\in\mathcal{U}$ be a minimizer of Problem 2 and set $m=G^{\prime}(H(x,Du))$. Then, there is a constant $C>0$, independent of $m$ and $u$, such that
$\begin{split}m\leq C\left|Du\right|^{\frac{\alpha}{\beta}}+C.\end{split}$
Proof.
Since the pair $(m,u)$ is a solution to Problem 1, we have that
$\begin{split}m=G^{\prime}(g(m))=G^{\prime}(H(x,Du)).\end{split}$
Subsequently, we have
$\begin{split}m=G^{\prime}(H(x,Du))&\leq C|H(x,Du)|^{\frac{1}{\beta}}+C\\ &\leq C|Du|^{\frac{\alpha}{\beta}}+C,\end{split}$
using the assumptions on $H$ and $G$. ∎
From this estimate, we obtain the following result.
Corollary 4.4.
Under Assumptions 1-4, and 5’ let $u$ solve Problem 2 and $m=G^{\prime}(H(x,Du))$. Then, $m\in L^{\beta+1}(\Omega)$, $g(m)\in L^{\frac{\beta+1}{\beta}}(\Omega)$, and $mD_{p}H(x,Du)\in\left(L^{\gamma^{\prime}}(\Omega)\right)^{d}$.
Proof.
Using the inequality $(a+b)^{r}\leq 2^{r}(a^{r}+b^{r})$, which holds for $a,b,r\in\mathbb{R}_{0}^{+}$, we have
$\begin{split}\int_{\Omega}m^{\beta+1}{\,\,\rm d}x&\leq C\int_{\Omega}|Du|^{\frac{(\beta+1)\alpha}{\beta}}{\,\,\rm d}x+C<\infty\end{split}$
since $\frac{(\beta+1)\alpha}{\beta}=\gamma$.
By Assumption 5, for $g(m)$, we have
$\begin{split}\int_{\Omega}g^{\frac{\beta+1}{\beta}}(m){\,\,\rm d}x\leq C\int_{\Omega}m^{\beta+1}{\,\,\rm d}x+C\leq C<\infty.\end{split}$
Now, set $\delta=\frac{\gamma-1}{\alpha-1}$ and $\delta^{\prime}=\frac{\gamma-1}{\gamma-\alpha}$, and note that $11/\delta+1/\delta^{\prime}=$ 1. Then,
$\begin{split}\gamma^{\prime}\delta(\alpha-1)&=\frac{\gamma}{\gamma-1}\,\frac{\gamma-1}{\alpha-1}\,(\alpha-1)=\gamma\quad\text{and}\quad\gamma^{\prime}\delta^{\prime}=\frac{\gamma}{\gamma-1}\frac{\gamma-1}{\gamma-\alpha}=\beta+1.\end{split}$
Due to Hölder’s inequality, we also have
$\begin{split}\int_{\Omega}\left|mD_{p}H(x,Du)\right|^{\gamma^{\prime}}&\leq\left[\int_{\Omega}m^{\gamma^{\prime}\delta^{\prime}}\right]^{\frac{1}{\delta^{\prime}}}\left[\int_{\Omega}\left|D_{p}H(x,Du)\right|^{\gamma^{\prime}\delta}\right]^{\frac{1}{\delta}}.\end{split}$
Using our assumptions on $H$ and the inequality $(a+b)^{r}\leq 2^{r}(a^{r}+b^{r})$, we further obtain
$\begin{split}\int_{\Omega}\left|mD_{p}H(x,Du)\right|^{\gamma^{\prime}}&\leq\left[\int_{\Omega}m^{\beta+1}\right]^{\frac{1}{\delta^{\prime}}}\left[\int_{\Omega}C(\left|Du\right|^{\alpha-1}+1)^{\gamma^{\prime}\delta}\right]^{\frac{1}{\delta}}\\ &\leq||m||_{\beta+1}^{\gamma^{\prime}}\left[\int_{\Omega}C2^{\gamma^{\prime}\delta}(\left|Du\right|^{(\alpha-1)\gamma^{\prime}\delta}+1)\right]^{\frac{1}{\delta}}\\ &=\widetilde{C}||m||_{\beta+1}^{\gamma^{\prime}}\left[||Du||_{\gamma}^{\gamma^{\prime}(\alpha-1)}+1\right]<C,\end{split}$
since $m\in L^{\beta+1}$ and $Du\in\left(L^{\gamma}(\Omega)\right)^{d}$. ∎
This leads to the following estimates.
Proposition 4.5.
Let $u$ solve Problem 2 and $m=G^{\prime}(H(x,Du))$. Then,
$\begin{split}\int_{\Omega}mg(m){\,\,\rm d}x<C,\qquad\int_{\Omega}\dfrac{(m+1)|Du|^{\alpha}}{C}<C.\end{split}$
Proof.
The first estimate is a direct consequence of Corollary 4.4.
For the second estimate, we recall that $m=G^{\prime}(H(x,Du))$, and $g(\cdot)$ is an inverse function of $G^{\prime}(\cdot)$, we have $H(x,Du)=g(m)$, and consider an estimate
$\begin{split}\int_{\Omega}(m+1)H(x,Du){\,\,\rm d}x=\int_{\Omega}(m+1)g(m){\,\,\rm d}x.\end{split}$
From Assumptions 5 and 4, we obtain
$\begin{split}\int_{\Omega}\dfrac{(m+1)|Du|^{\alpha}}{C}{\,\,\rm d}x\leq\int_{\Omega}(m+1)(g(m)+C){\,\,\rm d}x\leq\int_{\Omega}Cmg(m)+C<C.\end{split}$
5 The Correspondence Result
This section establishes the correspondence result stated in Theorem 1.2. For clarity, the proof is divided into two major parts, each addressed in a dedicated subsection. Recall that we are working under Assumption 5 and Assumption 5’, which are consistent due to Lemma 2.3.
5.1 From the variational formulation to the MFG
In this subsection, we rigorously establish that the MFG system (1.1) arises as the Euler–Lagrange equation of Problem 2 and that the boundary conditions (1.2) are naturally induced.
Proof of Theorem 1.2 (I).
Let $u\in\mathcal{U}$ be a minimizer of (1.9) and set
$m(x):=G^{\prime}(H(x,Du)).$ (5.1)
We now, calculate the Gateaux derivative using four different perturbation functions $v\in C^{\infty}(\mkern 1.0mu\overline{\mkern-1.0mu\Omega\mkern-1.0mu}\mkern 1.0mu)$.
  1. 1.
    Interior equations: Consider a function $v\in C_{c}^{\infty}(\Omega)$. Define a map $i(\varepsilon)=\mathcal{I}[u+\varepsilon v]$ for all $\varepsilon\in\mathbb{R}$. We see that $i(0)\leq i(\varepsilon)$, hence, $i^{\prime}(0)=0$. In particular, due to Lemma 2.4 and the dominated convergence theorem, we have
    $\begin{split}0&=\frac{d}{d\varepsilon}\int_{\Omega}G\left(H(x,Du+\varepsilon Dv)\right){\,\,\rm d}x\Big{|}_{\varepsilon=0}\\ &=\int_{\Omega}G^{\prime}\left(H(x,Du)\right)D_{p}H(x,Du)\cdot Dv{\,\,\rm d}x\\ \end{split}$
    Since $v\in C_{c}^{\infty}(\Omega)$ is arbitrary, we have $\operatorname*{div}(mD_{p}H(x,Du))=0$ in the sense of distributions. Accordingly, we obtain
    $\begin{cases}G^{\prime}(H(x,Du))=m&\text{ for all }x\in\Omega,\\ -\operatorname*{div}(mD_{p}H(x,Du))=0&\text{ for all }x\in\Omega,\end{cases}$ (5.2)
    which is the PDE system governing the interior structure of the MFG.
  2. 2.
    The Neumann boundary condition: Consider a function $v\in C^{\infty}(\mkern 1.0mu\overline{\mkern-1.0mu\Omega\mkern-1.0mu}\mkern 1.0mu)$ such that $v\equiv 0$ on $\Gamma_{D}$. Define a map $i(\varepsilon)=\mathcal{I}[u+\varepsilon v]$ for all $\varepsilon\in\mathbb{R}$. We see that $i(0)\leq i(\varepsilon)$, hence, $i^{\prime}(0)=0$. In particular,
    $\begin{split}0&=\frac{d}{d\varepsilon}\left(\int_{\Omega}G\left(H(x,Du+\varepsilon Dv)\right){\,\,\rm d}x-\int_{\Gamma_{N}}j(u+\varepsilon v){\,\,\rm d}s\right)\Big{|}_{\varepsilon=0}\\ &=\int_{\Omega}G^{\prime}\left(H(x,Du)\right)D_{p}H(x,Du)\cdot Dv{\,\,\rm d}x-\int_{\Gamma_{N}}jv{\,\,\rm d}s\\ &=\int_{\Omega}m\,D_{p}H(x,Du)\cdot Dv{\,\,\rm d}x-\int_{\Gamma_{N}}jv{\,\,\rm d}s.\end{split}$ (5.3)
    The previous steps ensures that $0=\operatorname*{div}(mD_{p}H(x,Du))\in L^{\gamma^{\prime}}(\Gamma)$. Also, because $u\in W^{1,\gamma}$ and $m\in L^{\beta+1}$, using Assumption 4, we get $mD_{p}H(x,Du)\in L^{\gamma^{\prime}}$. Thus, $mD_{p}H(x,Du)\in W^{\gamma^{\prime}}(\operatorname*{div};\Omega)$. Accordingly, we can invoke Theorem 2.8, and establish that
    $\int_{\Gamma_{N}}j\,v{\,\,\rm d}s=\int_{\Omega}m\,D_{p}H(x,Du)\cdot Dv{\,\,\rm d}x=\int_{\Gamma_{N}}v\,mD_{p}H(x,Du)\cdot\nu{\,\,\rm d}s.$
    Since $v$ is arbitrary, we obtain
    $mD_{p}H(x,Du)\cdot\nu=j\qquad\text{ for all }x\in\Gamma_{N}.$ (5.4)
  3. 3.
    The no-entry condition: Consider a function $v\in C^{\infty}(\mkern 1.0mu\overline{\mkern-1.0mu\Omega\mkern-1.0mu}\mkern 1.0mu)$ such that $v(x)\leq 0$ for all $x\in\Gamma_{D}$. Define a map $i(\varepsilon)=\mathcal{I}[u+\varepsilon v]$ for all $\varepsilon\geq 0$. Note that $u+\varepsilon v\in\mathcal{U}$ for all $\varepsilon\geq 0$; hence, $i(\varepsilon)\geq i(0)$. Using (5.2) and (5.4) above, we subsequently have
    $\begin{split}0\leq i^{\prime}(0)&=\int_{\Omega}m\,D_{p}H(x,Du)\cdot Dv{\,\,\rm d}x-\int_{\Gamma_{N}}j\,v{\,\,\rm d}s\\ &=\int_{\Gamma}v\,mD_{p}H(x,Du)\cdot\nu{\,\,\rm d}s-\int_{\Gamma_{N}}j\,v{\,\,\rm d}s=\int_{\Gamma_{D}}v\,mD_{p}H(x,Du)\cdot\nu{\,\,\rm d}s.\end{split}$
    Since $v|_{\Gamma}\leq 0$ is arbitrary, we obtain
    $\begin{split}mD_{p}H(x,Du)\cdot\nu\leq 0\qquad\text{on }\Gamma_{D}.\end{split}$ (5.5)
  4. 4.
    The contact-set condition: Consider an arbitrary function $v\in C^{\infty}(\mkern 1.0mu\overline{\mkern-1.0mu\Omega\mkern-1.0mu}\mkern 1.0mu)$. Define a map $i(\varepsilon):=\mathcal{I}[u+\varepsilon(\psi-u)v]$ for all $\varepsilon\in(-1,1)$ small enough so that $\sup_{\Gamma_{D}}|\varepsilon v|<1$. Note that,
    $v(\psi-u)\in W^{1,\gamma}(\Omega)\quad\text{ and }\quad u+\varepsilon v(\psi-u)\leq\psi\enskip\text{ on }\Gamma_{D};$
    hence, $u+\varepsilon v(\psi-u)\in\mathcal{U}$. Subsequently, $i^{\prime}(0)=0$ and, similar to before,
    $0=\int_{\Gamma_{D}}v(\psi-u)\,mD_{p}H(x,Du)\cdot\nu{\,\,\rm d}s.$
    Since $v$ is arbitrary, we obtain
    $\begin{split}(\psi-u)mD_{p}H(x,Du)\cdot\nu{\,\,\rm d}s=0\qquad\text{on }\Gamma_{D}.\end{split}$ (5.6)
Combining (5.2) with (5.4)–(5.6) gives us the MFG (1.1)-(1.2), implying that the MFG System in Problem 1 is the Euler-Lagrange equation for the variational formulation in Problem 2. Consequently, any minimizer of Problem 2 is a solution to Problem 1. $\square$.
5.2 From the MFG to the variational formulation
What we have established above in Subsection 5.1 is the first part of Theorem 1.2. Namely, we proved that a minimizer for Problem 2 produces a weak solution to Problem 1. In this subsection, we prove the converse, which states that a weak solution to Problem 1 produces a minimizer for Problem 2.
Proof of Theorem 1.2 (II).
Let $(m,u)$ be a weak solution to Problem 1 in the sense of Definition 2.11. We will show that for all $w\in\mathcal{U}$,
$\begin{split}\mathcal{I}[u]\leq\mathcal{I}[w].\end{split}$
Indeed, due to the convexity of $G$ and $H$, we have
$\mathcal{I}[w]-\mathcal{I}[u]=$ $\int_{\Omega}G(H(x,Dw))-G(H(x,Du)){\,\,\rm d}x-\int_{\Gamma_{N}}j(w-u){\,\,\rm d}s$
$\geq$ $\int_{\Omega}G^{\prime}(H(x,Du))D_{p}H(x,Du)D(w-u){\,\,\rm d}x-\int_{\Gamma_{N}}j(w-u){\,\,\rm d}s$
$=$ $\int_{\Omega}mD_{p}H(x,Du)D(w-u){\,\,\rm d}x-\int_{\Gamma_{N}}j(w-u){\,\,\rm d}s,$
where we used that $G^{\prime}(H(x,Du))=G^{\prime}(g(m))=m$. Now, by our definition of weak solutions (Definition 2.11) and the normal trace theorem (Theorem 2.8), we obtain
$\begin{split}\mathcal{I}[w]-\mathcal{I}[u]\geq&\int_{\Omega}mD_{p}H(x,Du)D(w-u){\,\,\rm d}x-\int_{\Gamma_{N}}j(w-u){\,\,\rm d}s\\ =&\int_{\Gamma}(w-u)\,mD_{p}H(x,Du)\cdot\nu{\,\,\rm d}s-\int_{\Gamma_{N}}j(w-u){\,\,\rm d}s\\ =&\int_{\Gamma_{D}}(w-u)\,mD_{p}H(x,Du)\cdot\nu{\,\,\rm d}s.\end{split}$
Finally, due to our assumptions, $(w-\psi)\,mD_{p}H(x,Du)\cdot\nu\geq 0$ and $(\psi-u)\,mD_{p}H(x,Du)\cdot\nu=0$ on $\Gamma_{D}$, giving us that
$\mathcal{I}[w]-\mathcal{I}[u]\geq\int_{\Gamma_{D}}[(w-\psi)+(\psi-u)]\,mD_{p}H(x,Du)\cdot\nu{\,\,\rm d}s\geq 0.$
This concludes the proof. $\square$
6 Monotonicity and the uniqueness of solution
In this section, we establish uniqueness results for the MFG system by employing the monotonicity techniques developed in [LL07]. We begin with a discussion on the related monotone MFG operator, which was introduced in [FGT19]. Subsequently, we demonstrate the uniqueness of the density function $m$, while the uniqueness of the gradient $Du$ is guaranteed in the region where the density remains strictly positive, i.e., where $m>0$.
6.1 The MFG operator
Set $\mathcal{X}:=L^{\beta+1}(\Omega)\times W^{1,\gamma}(\Omega)$ and let $\mathcal{X}^{\prime}$ be its dual. Let $\mathcal{Y}\subset\mathcal{X}$ be the family of all functions $(m,u)\in\mathcal{X}$ that satisfy the conditions:
  1. i.
    The density $m$ is nonnegative.
  2. ii.
    The Dirichlet boundary condition in (1.2) is satisfied in the trace sense.
  3. iii.
    The divergence of the flow, $-\operatorname*{div}(mD_{p}H(x,Du))$, is in $L^{\gamma^{\prime}}(\Omega)$.
  4. iv.
    The Neumann boundary conditions in (1.2) are satisfied in the Neumann trace sense.
The MFG operator $A:\mathcal{Y}\to\mathcal{X}^{\prime}$ we investigate in this section is given by
$A\begin{bmatrix}m\\[1.00006pt] u\end{bmatrix}=\begin{bmatrix}-H(x,Du)+g(m)\\[2.5pt] -\operatorname*{div}(mD_{p}H(x,Du))\end{bmatrix},$ (6.1)
which is was introduced in [FGT19] inspired by the uniqueness proof in [LL07]. The action of this operator is given by the pairing
$\begin{split}\Big{\langle}A\begin{bmatrix}m\\[1.00006pt] u\end{bmatrix}\,,\,\begin{bmatrix}\mu\\[1.00006pt] v\end{bmatrix}\Big{\rangle}=\int_{\Omega}(-H(x,Du)+g(m))\,\mu{\,\,\rm d}x&-\int_{\Omega}\operatorname*{div}(mD_{p}H(x,Du))\,v{\,\,\rm d}x\\ =\int_{\Omega}(-H(x,Du)+g(m))\,\mu{\,\,\rm d}x&+\int_{\Omega}mD_{p}H(x,Du)\cdot Dv{\,\,\rm d}x\\ &\phantom{44444}-\int_{\Gamma}v\,\,mD_{p}H(x,Du)\cdot\nu{\,\,\rm d}s.\end{split}$ (6.2)
This subsection’s main objective is to establish the MFG operator’s monotonicity. This property, which is defined as follows, allows us to prove the MFG’s uniqueness results.
Definition 6.1.
Let $\mathcal{X}$ be a Banach space and $\mathcal{X}^{\prime}$ its dual. An operator $A:\mathcal{Y}\subset\mathcal{X}\to\mathcal{X}^{\prime}$ is monotone if and only if, for all $a,b\in\mathcal{Y}$, it satisfies
$\begin{split}\left\langle A(a)-A(b)\,\,,\,\,a-b\right\rangle\geq 0.\end{split}$
Remark 6.2.
The well-definedness of the pairing in (6.2) relies on the integrability of the constituent terms. While Assumptions 15 ensure that
$-H(x,Du)+g(m)\in L^{(\beta+1)/\beta}(\Omega)\quad\text{and}\quad mD_{p}H(x,Du)\in L^{\gamma^{\prime}}(\Omega;\mathbb{R}^{d})$
for all pairs $(m,u)\in L^{\beta+1}(\Omega)\times W^{1,\gamma}(\Omega)$, imposing the additional constraint
$\operatorname*{div}(mD_{p}H(x,Du))\in L^{\gamma^{\prime}}(\Omega)$
allows us to apply the normal trace theorem (Theorem 2.8) and guarantees that
$mD_{p}H(x,Du)\cdot\nu\in W^{1-1/\gamma^{\prime},\gamma^{\prime}}(\Gamma).$
6.2 The monotonicity of the MFG operator
The following result establishes the monotonicity of the mean-field game operator.
Proposition 6.3.
The operator $A:\mathcal{Y}\to\mathcal{X}^{\prime}$, as defined in (6.1), is monotone on $\mathcal{Y}$.
Proof.
Let $(m,u)$ and $(\mu,v)$ be two pairs in $\mathcal{Y}$. We have
$\begin{split}\Big{\langle}A\begin{bmatrix}m\\[1.00006pt] u\end{bmatrix}-A\begin{bmatrix}\mu\\[1.00006pt] v\end{bmatrix}\,&,\,\begin{bmatrix}m-\mu\\[1.00006pt] u-v\end{bmatrix}\Big{\rangle}\\ =&\int_{\Omega}\left[g(m)-g(\mu)\right](m-\mu){\,\,\rm d}x\\ &\quad+\int_{\Omega}\left[-H(x,Du)+H(x,Dv)\right](m-\mu){\,\,\rm d}x\\ &\quad+\int_{\Omega}\left[mD_{p}H(x,Du)-\mu D_{p}H(x,Dv)\right](Du-Dv){\,\,\rm d}x\\ &\quad+\int_{\Gamma}\left[-mD_{p}H(x,Du)+\mu D_{p}H(x,Dv)\right]\cdot\nu(u-v){\,\,\rm d}s.\end{split}$ (6.3)
Now, we notice that since $(m,u)$ and $(\mu,v)$ satisfy the boundary conditions (1.2), we have, on the one hand,
$mD_{p}H(x,Du)\cdot\nu=\mu D_{p}H(x,Dv)\cdot\nu=j\qquad\text{on }\Gamma_{N}.$ (6.4)
On the other hand, on $\Gamma_{D}$, we have
$\begin{split}-(u-\psi)mD_{p}H(x,Du)\cdot\nu&=0,\qquad(\psi-v)\mu D_{p}H(x,Dv)\cdot\nu=0,\\ -(\psi-v)mD_{p}H(x,Du)\cdot\nu&\geq 0,\qquad(u-\psi)\mu D_{p}H(x,Dv)\cdot\nu\geq 0.\end{split}$ (6.5)
Combining (6.4) and (6.5), we obtain that
$\begin{split}I_{1}:=\int_{\Gamma}(u-v)\left[-mD_{p}H(x,Du)+\mu D_{p}H(x,Dv)\right]\cdot\nu{\,\,\rm d}s\geq 0.\end{split}$ (6.6)
Moreover, because $H(x,p)$ is convex in $p$, we have
$\begin{split}H(x,Du)-H(x,Dv)&\geq(Du-Dv)\cdot D_{p}H(x,Dv),\end{split}$
and
$\begin{split}H(x,Dv)-H(x,Du)&\geq(Dv-Du)\cdot D_{p}H(x,Du).\end{split}$
Multiplying the first inequality by $\mu$ and the second by $m$ and summing the two equations, we obtain
$\begin{split}I_{2}:=\int_{\Omega}(m-\mu)&\left(H(x,Dv)-H(x,Du)\right)\\ +&\int_{\Omega}\left(mD_{p}H(x,Du)-\mu D_{p}H(x,Dv)\right)\cdot(Du-Dv){\,\,\rm d}x\geq 0.\end{split}$ (6.7)
Finally, since $g$ is monotone, we have that
$I_{3}:=\int_{\Omega}(g(m)-g(\mu))(m-\mu){\,\,\rm d}x\geq 0.$ (6.8)
Due to inequalities (6.6), (6.7), and (6.8), inequality (6.3) becomes
$\begin{split}\Big{\langle}A\begin{bmatrix}m\\[1.00006pt] u\end{bmatrix}-A\begin{bmatrix}\mu\\[1.00006pt] v\end{bmatrix}\,,\,\begin{bmatrix}m-\mu\\[1.00006pt] u-v\end{bmatrix}\Big{\rangle}=I_{1}+I_{2}+I_{3}\geq 0.\end{split}$
This concludes the proof. ∎
6.3 On the Uniqueness of Solution to the MFG
Here, we use the previously proven monotonicity to establish that $Du$ is unique in the region where the density function $m$ is strictly positive, implying that the density function $m$ itself is also unique in $\Omega$.
Proof of Theorem 1.3.
The proof utilizes the Lasry-Lions monotonicity method [LL06b].
Using the notation in the proof of Proposition 6.3, we see that
$\begin{split}A\begin{bmatrix}m\\[1.00006pt] u\end{bmatrix}=A\begin{bmatrix}\mu\\[1.00006pt] v\end{bmatrix}=\begin{bmatrix}0\\ 0\end{bmatrix}.\end{split}$
Therefore, due to inequalities (6.6), (6.7), and (6.8), we have $I_{1}=I_{2}=I_{3}=0$. By the inequalities in (6.5), we see that $I_{1}=0$ implies that
$\begin{split}(\psi-u)\eta D_{p}H(x,Dv)\cdot\nu=0\quad\text{and}\quad-(\psi-v)mD_{p}H(x,Du)\cdot\nu=0.\end{split}$
This establishes 1. Now, since $g$ is strictly monotone, we have that
$\begin{split}(g(m)-g(\eta))(m-\eta)>0\end{split}$
whenever $\eta\neq m$. However, because $I_{3}=0$, we have that $(g(m)-g(\eta))(m-\eta)=0$ almost everywhere in $\Omega$. Hence, $\eta=m$ almost everywhere in $\Omega$. This establishes 2. For 3, we use the strict convexity assumption on $H(x,p)$ in the variable $p$. Indeed, given that $I_{2}=0$, we see that whenever $m=\eta\neq 0$, then
$\begin{split}H(x,Du)-H(x,Dv)=(Du-Dv)\cdot D_{p}H(x,Dv).\end{split}$
However, this is only possible when $Du=Dv$. This ends the proof. $\square$
6.4 On the uniqueness of minimizer
Given compatible coupling term $g$ and mapping $G$, the inherent correspondence established in Theorem 1.2 between Problem 1 and Problem 2 naturally induces a correspondence between results in the MFG setting and the variational formulation setting. In the variational formulation setting, this correspondence gives rise to the following two results that serve as the counterpart to the uniqueness result of Theorem 1.3.
Corollary 6.4.
Under Assumptions 14 and Assumption 5’, let $u,\,v\in\mathcal{U}$ be two minimizers of Problem 2. The following must hold:
  1. 1.
    For almost all $x\in\Gamma_{D}$, we have
    $(\psi-u)G^{\prime}(H(x,Dv))D_{p}H(x,Dv)\cdot\nu=(\psi-v)G^{\prime}(H(x,Du))D_{p}H(x,Du)\cdot\nu=0.$
  2. 2.
    For almost all $x\in\Omega$,
    $G^{\prime}\bigl{(}H(x,Du(x))\bigr{)}=G^{\prime}\bigl{(}H(x,Dv(x))\bigr{)}.$
  3. 3.
    For almost all $x\in\Omega$,
    $G^{\prime}\bigl{(}H(x,Du(x))\bigr{)}\neq 0\implies Du(x)=Dv(x).$
Proof.
The claim follows directly from Theorem 1.2 and Theorem 1.3. ∎
As a subsequence of Corollary 6.4 above, we obtain the following corollary related to boundary behavior.
Corollary 6.5.
Under Assumptions 1-4 and Assumption 5’, let $u,\,v\in\mathcal{U}$ be two minimizers of Problem 2. The following must hold:
  1. 1.
    For almost all $x\in\Omega$,
    $G\bigl{(}H(x,Du(x))\bigr{)}=G\bigl{(}H(x,Dv(x))\bigr{)}.$
  2. 2.
    For almost all $x\in\Omega$, if
    $\int_{\Gamma_{N}}j\,v{\,\,\rm d}s=\int_{\Gamma_{N}}ju{\,\,\rm d}s.$
Proof.
  1. 1.
    Due Corollary 6.4.3, we immediately obtain that
    $G^{\prime}(H(x,Du))\neq 0\implies G(H(x,Du))=G(H(x,Dv)).$
    On the other hand, let
    $G_{0}=\inf_{z\in\mathbb{R}}G(z).$
    Since $G$ is increasing, whenever $G^{\prime}(z)=0$, we have that $G(z)=G_{0}$. Therefore,
    $G^{\prime}(H(x,Du))=0\implies G(H(x,Du))=G(H(x,Du))=G_{0}.$
    This establishes the first claim.
  2. 2.
    The second claim follows from the first. Because $\mathcal{I}[u]=\mathcal{I}[v]$, we have
    $\begin{split}\mathcal{I}[u]&=\int_{\Omega}G\left(H(x,Du)\right){\,\,\rm d}x-\int_{\Gamma_{N}}ju{\,\,\rm d}s\\ &=\int_{\Omega}G\left(H(x,Dv)\right){\,\,\rm d}x-\int_{\Gamma_{N}}jv{\,\,\rm d}s+\int_{\Gamma_{N}}jv-\int_{\Gamma_{N}}ju\\ &=\mathcal{I}[v]+\int_{\Gamma_{N}}jv-\int_{\Gamma_{N}}ju.\end{split}$
    Therefore,
    $\int_{\Gamma_{N}}jv=\int_{\Gamma_{N}}ju.$
This concludes the proof. ∎
Appendix
Appendix A The Neumann Trace and Neumann Boundary Condition
In this appendix, we address the weak formulation of the Neumann boundary condition for our MFG system under low-regularity assumptions, using the normal (Neumann) trace theorem. A rigorous proof of the well-definedness of this operator is provided in [Gal11] for the Banach-space $\tilde{H}_{p}(\Omega)$, corresponding to $W^{p}(\operatorname*{div};\Omega)$ in our notation. A more accessible outline of the theorem in the Hilbert-space setting $H(\operatorname*{div};\Omega)$ appears in [Tar07]. For completeness, we present the proof of this result, drawing on the techniques of [Tar07], [Gal11], [DiB16], and [Eva10].
We note that, to maintain the coherence of this appendix, we use the standard exponent $p$ for our $L^{p}$–spaces and denote its convex conjugate by $q$. So, in what follows, $\frac{1}{p}+\frac{1}{q}=1$. For the main body of this paper, we substitute $p=\gamma^{\prime}$ and $q=\gamma$. We use the term Neumann trace to avoid ambiguity when discussing the usual boundary trace in the same context.
A.1 The weak divergence
Let $\Omega\subset\mathbb{R}^{d}$ be an open, bounded domain with smooth boundary $\Gamma=\partial\Omega$. Given sufficiently regular vector field $\mathbf{u}$ and function $\phi$, the divergence theorem can be expressed as the integration-by-parts identity
$\int_{\Omega}\phi\,\operatorname*{div}(\mathbf{u}){\,\,\rm d}x=-\int_{\Omega}\mathbf{u}\cdot D\phi{\,\,\rm d}x+\int_{\Gamma}\phi\enskip\mathbf{u}\cdot\nu{\,\,\rm d}s.$ (a.1)
This identity constitutes the foundational structure for the weak formulation of Neumann boundary conditions in divergence-form PDE. Our goal is to provide sufficient conditions for the Neumann trace $\mathbf{u}\cdot\nu$ to be well-defined and to extend this identity to a larger class of vector fields whose divergence may only be weakly defined in the following sense.
Definition A.1.
The weak divergence of a vector-field $\mathbf{u}\in L^{1}(\Omega;\mathbb{R}^{d})$ is the linear functional $\operatorname*{div}(\mathbf{u})\in L^{1}_{loc}(\Omega)$ that satisfies
$\langle\operatorname*{div}(\mathbf{u})\,,\phi\rangle:=\int_{\Omega}\operatorname*{div}(\mathbf{u})\phi=-\int_{\Omega}\mathbf{u}\cdot D\phi{\,\,\rm d}x$ (a.2)
for every $\phi\in C_{c}^{\infty}(\Omega)$.
To extend this definition to (a.1), which incorporates boundary traces, we need to use a class of test functions larger than $\mathcal{D}(\Omega)=C^{\infty}_{c}(\Omega)$. In particular, we use the following family of test functions.
Definition A.2.
We define the family of test functions $\mathfrak{D}(\mkern 1.0mu\overline{\mkern-1.0mu\Omega\mkern-1.0mu}\mkern 1.0mu)$ by
$\{\,\phi|_{\mkern 1.0mu\overline{\mkern-1.0mu\Omega\mkern-1.0mu}\mkern 1.0mu}\,\colon\,\phi\in C^{\infty}(\mathbb{R}^{d})\,\},$
which is the set of restrictions of $C^{\infty}$–functions that are defined on the entire Euclidean space $\mathbb{R}^{d}$.
Remark A.3.
We denote this family of test functions $\mathfrak{D}(\mkern 1.0mu\overline{\mkern-1.0mu\Omega\mkern-1.0mu}\mkern 1.0mu)$ to avoid ambiguity with commonly used family of test functions $\mathcal{D}(\Omega)=C^{\infty}_{c}(\Omega)$.
A.2 Divergence Sobolev space
For the Neumann trace to be well-defined in the framework of weak solutions, we need to choose an appropriate function space and a domain $\Omega$ whose boundary enjoys an appropriate level of regularity. The space we are considering in this section is the following divergence Sobolev space.
Definition A.4.
Let $\Omega\subset\mathbb{R}^{d}$ be an open, bounded domain, and suppose that $1\leq p<\infty$. The divergence Sobolev space $W^{p}(\operatorname*{div};\Omega)$ is the space of vector fields
$\begin{split}W^{p}(\operatorname*{div};\Omega):=\{\mathbf{w}\in L^{p}(\Omega;\mathbb{R}^{d})\,:\,\operatorname*{div}(\mathbf{w})\in L^{p}(\Omega)\},\end{split}$
where $\operatorname*{div}(\mathbf{w})$ is defined weakly according to Definition A.1. This space is equipped with the natural norm
$\begin{split}\|\mathbf{w}\|_{W^{p}(\operatorname*{div};\Omega)}:=\Big{(}\|\mathbf{w}\|_{L^{p}(\Omega;\mathbb{R}^{d})}^{p}+\|\operatorname*{div}(\mathbf{w})\|_{L^{p}(\Omega)}^{p}\Big{)}^{1/p}.\end{split}$
Remark A.5.
We use this name divergence Sobolev space consistently with common divergence-measure fields known in the literature [CF01].
The regularity condition we require the domain $\Omega$ to have on its boundary is the following variant of the segment condition (see [Eva10], [DiB16]).
Definition A.6.
A bounded domain $\Omega\subset\mathbb{R}^{d}$ satisfies the segment condition on its boundary $\Gamma=\partial\Omega$ if there are two finite collections
$\{x_{n}\}_{n=1}^{N}\subset\Gamma\quad\text{and}\quad\{y_{n}\}_{n=1}^{N}\subset\mathbb{R}^{d}\backslash\{0\}$
and two positive constants $r$ and $t^{*}$ such that
  1. i.
    the collection $\mathcal{B}_{N}=\{B_{r}(x_{n})\}_{n=1}^{N}$ of open balls covers $\Gamma$, and
  2. ii.
    for $n=1,...,N$, we have
    $\begin{split}B_{t}(x+ty_{n})\subset\Omega\end{split}$
    for all $x\in B_{r}(x_{n})\cap\mkern 1.0mu\overline{\mkern-1.0mu\Omega\mkern-1.0mu}\mkern 1.0mu$ and $t\in(0,t^{*})$.
Remark A.7.
This condition holds when $\Gamma$ is $C^{1}$ [Eva10].
Lemma A.8.
Let $\Omega\subset\mathbb{R}^{d}$ be an open, bounded domain. The space $W^{p}(\operatorname*{div};\Omega)$ is a Banach space.
Proof.
We only prove the completeness property. The remaining normed–space properties are trivial.
  1. 1.
    Consider an arbitrary Cauchy sequence in $\left(\mathbf{u}_{n}\right)_{n\geq 1}$ in $W^{p}(\operatorname*{div};\Omega)$. Then, the sequence $\left(\mathbf{u}_{n}\right)_{n\geq 1}$ and $\left(\operatorname*{div}(\mathbf{u}_{n})\right)_{n\geq 1}$ are Cauchy sequences in $L^{p}(\Omega;\mathbb{R}^{d})$ and $L^{p}(\Omega)$, respectively. By the completeness of $L^{p}$–spaces, there is a vector field $\mathbf{u}\in L^{p}(\Omega;\mathbb{R}^{d})$ and a function $v\in L^{p}(\Omega)$ such that
    $\mathbf{u}_{m}\to\mathbf{u}\quad\text{and}\quad\operatorname*{div}(\mathbf{u}_{m})\to v\in L^{p}(\Omega).$
  2. 2.
    Note that
    $\begin{split}\int_{\Omega}\operatorname*{div}(\mathbf{u})\phi=-\int_{\Omega}\mathbf{u}\cdot D\phi&=\lim_{n\to\infty}-\int_{\Omega}\mathbf{u}_{n}\cdot D\phi\\ &=\lim_{n\to\infty}\int_{\Omega}\operatorname*{div}(\mathbf{u}_{n})\phi=\int_{\Omega}v\phi.\end{split}$
    for all $\phi\in C^{\infty}_{c}(\Omega)$. Due to the density of $\phi\in C^{\infty}_{c}(\Omega)$ in $L^{p^{\prime}}(\Omega)$, we have that
    $\|\operatorname*{div}(\mathbf{u})-v\|_{L^{p}(\Omega)}=\sup_{\begin{subarray}{c}\phi\in C^{\infty}_{c}(\Omega)\\ \|\phi\|\leq 1\end{subarray}}\int_{\Omega}(\operatorname*{div}(\mathbf{u})-v)\phi=0.$
    This shows that $\operatorname*{div}(\mathbf{u})=v$ almost everywhere in $\Omega$.
This concludes the proof. ∎
Our goal in this section is to define a trace operator on $W^{p}(\operatorname*{div};\Omega)$ in a way that is consistent with the usual trace definition. We start by establishing the following result.
Proposition A.9.
[…]
Proof.
Let $\mathbf{u}=(u^{1},...,u^{d})\in W^{p}(\operatorname*{div};\Omega)$.
  1. 1.
    Let $B_{r}({x_{0}})\in\mathcal{B}_{N}$, where $\mathcal{B}_{N}$ is the cover from the segment condition, and let $y_{0}$ be the vector for which $B_{t}(x+ty_{0})\subset\Omega$ for all $t\in(0,t^{*})$ and $x\in\mkern 1.0mu\overline{\mkern-1.0mu\Omega\mkern-1.0mu}\mkern 1.0mu\cap B_{r}(x_{0})$.
  2. 2.
    For every $t\in(0,t^{*})$ and $x\in\mkern 1.0mu\overline{\mkern-1.0mu\Omega\mkern-1.0mu}\mkern 1.0mu\cap B_{r}(x_{0})$, define $\mathbf{u}_{t}$ to be
    $\mathbf{u}_{t}(x)=(\mathbf{u}^{1}_{t},...,\mathbf{u}^{d}_{t}):=\mathbf{u}(x+ty_{0}),$
    and define $\mathbf{v}_{t}$ to be its mollified version
    $\mathbf{v}_{t}:=\eta_{t}\ast\mathbf{u}_{t}(x),$
    where $\eta_{t}(x)=c_{t}\eta(x/t)$ is the standard mollifier.
  3. 3.
    First, note that
    $\|\operatorname*{div}\mathbf{v}_{t}-\operatorname*{div}\mathbf{u}\|_{L^{p}(B_{r}(x_{0}))}\leq\|\operatorname*{div}\mathbf{v}_{t}-\operatorname*{div}\mathbf{u}_{t}\|_{L^{p}(B_{r}(x_{0}))}+\|\operatorname*{div}\mathbf{u}_{t}-\operatorname*{div}\mathbf{u}\|_{L^{p}(B_{r}(x_{0}))}.$ (a.3)
    On the right-hand side, the second term vanishes in the limit due to the continuity of translations in $L^{p}$ spaces. The first term vanishes in the limit following the technique used for the classical weak derivative (see [Eva10]). Thus, as $t\to 0$, $\mathbf{v}_{t}\to\mathbf{u}$ strongly in $W^{p}(\operatorname*{div};\Omega)$.
  4. 4.
    Let $\Omega_{0}\subset\subset\Omega$ be an open subset, compactly contained in $\Omega$, such that
    $\Omega\subset\Omega_{0}\cup\;\bigcup_{n=1}^{N}B_{r}(x_{n}).$
    Denote by $\mathbf{v}^{n}_{t}$ the smooth approximation of $\mathbf{u}$ in $B_{r}(x_{n})$. Also, denote by $\mathbf{v}^{0}_{t}$ the smooth approximation of $\mathbf{u}$ in $\Omega_{0}$, which exists following the above procedure without the need for translation. Let $\{\xi_{n}\}_{n=0}^{N}$ be a smooth partition of unity corresponding to the open cover above, and define
    $\mathbf{v}_{t}=\sum_{n=0}^{N}\xi_{n}\mathbf{v}_{t}^{n}.$
    Since the summation is finite, and $\sum_{n=0}^{N}\xi_{n}=1$, we have that
    $\begin{split}\|\mathbf{v}_{t}-\mathbf{u}\|_{W^{p}(\operatorname*{div};\Omega)}&\leq\sum_{n=0}^{N}(\|\operatorname*{div}(\xi_{n}\mathbf{v}^{n}_{t})-\operatorname*{div}(\xi_{n}\mathbf{u})\|_{L^{p}(\Omega)}+\|\xi_{n}\mathbf{v}^{n}_{t}-\xi_{n}\mathbf{u}\|_{L^{p}(\Omega;\mathbb{R}^{d})})\\ &\leq C\|\mathbf{v}^{0}_{t}-\mathbf{u}\|_{W^{p}(\operatorname*{div};\Omega_{0})}+C\sum_{n=1}^{N}\|\mathbf{v}^{n}_{t}-\mathbf{u}\|_{W^{p}(\operatorname*{div};B_{r}(x_{n}))},\end{split}$
    which goes to 0 as $t\to 0$. ∎
Remark A.10.
We omitted the part on the convergence of $\|\xi_{n}\mathbf{v}^{n}_{t}-\xi_{n}\mathbf{u}\|_{L^{p}(\Omega;\mathbb{R}^{d})}$ since it is standard (see [Eva10] for example).
A.3 The Neumann Trace
Let $q$ denotes the convex (Hölder) conjugate of $p$; i.e., $\frac{1}{p}+\frac{1}{q}=1$. Let $\operatorname{\mathfrak{tr}}_{0}$ be the usual (Dirichlet) trace operator on the space $W^{1,q}(\Omega)$. We observe that, given a smooth vector-field $\mathbf{u}\in\left(\mathfrak{D}(\mkern 1.0mu\overline{\mkern-1.0mu\Omega\mkern-1.0mu}\mkern 1.0mu)\right)^{d}$ and a function $v\in W^{1,q}(\Omega)$, we have
$\begin{split}\int_{\Omega}\left[\mathbf{u}\cdot Dv+v\,\operatorname*{div}\mathbf{u}\right]\,{\,\,\rm d}x=\int_{\Gamma}\mathbf{u}\cdot\nu\enskip\operatorname{\mathfrak{tr}}_{0}v\,{\,\,\rm d}s,\end{split}$ (a.4)
where $\nu$ is the exterior normal to $\Gamma$. We aim to use the density result above to extend the Neumann trace $\mathbf{u}\cdot\nu$ to vector-fields $\mathbf{u}\in W^{p}(\operatorname*{div};\Omega)$ using this identity. In particular, we want to define $\mathbf{u}\cdot\nu:=\operatorname{\mathfrak{tr}}_{1}(\mathbf{u})$ as the $W^{p}(\operatorname*{div};\Omega)$–completion of a bounded linear map $\operatorname{\mathfrak{tr}}_{1}:\left(\mathfrak{D}(\mkern 1.0mu\overline{\mkern-1.0mu\Omega\mkern-1.0mu}\mkern 1.0mu)\right)^{d}\to W^{-1+\frac{1}{q},p}(\Gamma)$, where
$W^{-1+\frac{1}{q},p}(\Gamma):=\left(W^{1-\frac{1}{q},q}(\Gamma)\right)^{\prime}$
is the dual of the trace space of $W^{1,q}(\Omega)$.
We establish the validity of this definition via the extension of functions $v\in W^{1-\frac{1}{q},q}(\Gamma)$, defined on the boundary, to the interior of the domain $\Omega$ via an appropriate lifting operator, which is provided by the following lemma.
Lemma A.11.
Let $\Omega$ be a bounded, open, connected, $C^{1}$–domain. Then, there exists a lifting operator $l_{0}:W^{1-\frac{1}{q},q}(\Gamma)\to W^{1,q}(\Omega)$ such that
$\begin{split}\|l_{0}v\|_{W^{1,q}(\Omega)}\leq C\|v\|_{W^{1-\frac{1}{q},q}(\Gamma)}\end{split}$
for every $v\in C^{\infty}(\Gamma)$. The constant $C$ depends only on $\Omega$ and $q$.
Proof.
See [DiB16], Chapter 10, Theorem 17.1c and Theorem 18.2c. ∎
Equipped with the preceding technical tools, we now extend the identity (a.4) to an arbitrary $\mathbf{u}\in W^{p}(\operatorname*{div};\Omega)$.
Theorem A.12.
Let $\Omega$ be an open, bounded, $C^{1}$–domain. Then, there is a bounded linear operator
$\begin{split}\operatorname{\mathfrak{tr}}_{1}:W^{p}(\operatorname*{div};\Omega)\to W^{-1+\frac{1}{q},\,p}(\Gamma)\end{split}$
satisfying the following properties.
  1. (i)
    The operator $\operatorname{\mathfrak{tr}}_{1}$ is consistent with point-wise evaluation on the boundary for smooth function; that is,
    $\operatorname{\mathfrak{tr}}_{1}(\mathbf{u})=\mathbf{u}|_{\Gamma}\cdot\nu\quad\text{ for all }\,\mathbf{u}\in\left(\mathfrak{D}(\mkern 1.0mu\overline{\mkern-1.0mu\Omega\mkern-1.0mu}\mkern 1.0mu)\right)^{d}.$
  2. (ii)
    For each $\mathbf{u}\in W^{p}(\operatorname*{div};\Omega)$, we have that $\operatorname{\mathfrak{tr}}_{1}(\mathbf{u})\in W^{-1+\frac{1}{q},p}(\Gamma)$. That is, $\operatorname{\mathfrak{tr}}_{1}(\mathbf{u})$ is linear functional on $W^{1-\frac{1}{q},q}(\Gamma)$ bounded by
    $\|\operatorname{\mathfrak{tr}}_{1}(\mathbf{u})\|_{W^{-1+\frac{1}{q},p}(\Gamma)}\leq C\|\mathbf{u}\|_{W^{p}(\operatorname*{div};\Omega)}$ (a.5)
    for some constant $C$ independent of $\mathbf{u}$.
Proof.
  1. 1.
    For any smooth vector-field $\mathbf{u}\in\left(\mathfrak{D}(\mkern 1.0mu\overline{\mkern-1.0mu\Omega\mkern-1.0mu}\mkern 1.0mu)\right)^{d}$ (see Definition A.2), we define the Neumann trace of $\mathbf{u}$ in the usual sense:
    $\operatorname{\mathfrak{tr}}_{1}(\mathbf{u})=\mathbf{u}|_{\Gamma}\cdot\nu.$
  2. 2.
    Fix $v\in W^{1-\frac{1}{q},q}(\Gamma)$, and define the linear functional
    $\Lambda_{v}\mathbf{u}=\langle\operatorname{\mathfrak{tr}}_{1}(\mathbf{u}),v\rangle$
    for all $\mathbf{u}\in\left(\mathfrak{D}(\mkern 1.0mu\overline{\mkern-1.0mu\Omega\mkern-1.0mu}\mkern 1.0mu)\right)^{d}$. Now, using the lifting operator from Lemma A.11, we obtain
    $\begin{split}|\Lambda_{v}\mathbf{u}|=|\langle\operatorname{\mathfrak{tr}}_{1}(\mathbf{u}),v\rangle|&=\big{|}\int_{\Gamma}\operatorname{\mathfrak{tr}}_{1}(\mathbf{u})\,v{\,\,\rm d}s\big{|}\\[2.5pt] &=\big{|}\int_{\Omega}\mathbf{u}\cdot D(l_{0}v)+\operatorname*{div}(\mathbf{u})\enskip l_{0}v{\,\,\rm d}x\big{|}\\[2.5pt] &\leq\|\mathbf{u}\|_{W^{p}(\operatorname*{div};\Omega)}\enskip\|l_{0}v\|_{W^{1,q}(\Omega)}\\[2.5pt] &\leq C\|\mathbf{u}\|_{W^{p}(\operatorname*{div};\Omega)}\enskip\|v\|_{W^{1-\frac{1}{q},q}(\Gamma)}.\end{split}$ (a.6)
    Therefore, $\Lambda_{v}$ is a bounded linear functional for all $\mathbf{u}\in\left(\mathfrak{D}(\mkern 1.0mu\overline{\mkern-1.0mu\Omega\mkern-1.0mu}\mkern 1.0mu)\right)^{d}\cap W^{p}(\operatorname*{div};\Omega)$.
  3. 3.
[…]
    $\Lambda_{v}\mathbf{u}=\lim\limits_{n\to\infty}\Lambda_{v}\mathbf{u}_{n}$
    for some sequence $\{\mathbf{u}_{n}\}_{n=1}^{\infty}\subset\left(\mathfrak{D}(\mkern 1.0mu\overline{\mkern-1.0mu\Omega\mkern-1.0mu}\mkern 1.0mu)\right)^{d}$ such that $\|\mathbf{u}-\mathbf{u}_{n}\|_{W^{p}(\operatorname*{div};\Omega)}\to 0$ as $n\to\infty$. This limit also preserves the bound
    $|\Lambda_{v}\mathbf{u}|\leq C\|\mathbf{u}\|_{W^{p}(\operatorname*{div};\Omega)}\enskip\|v\|_{W^{1-\frac{1}{q},q}(\Gamma)}.$
  4. 4.
    Therefore, we simply define $\operatorname{\mathfrak{tr}}_{1}(\mathbf{u})$ for an arbitrary $\mathbf{u}\in W^{p}(\operatorname*{div};\Omega)$ according to the pairing
    $\langle\operatorname{\mathfrak{tr}}_{1}(\mathbf{u}),v\rangle=\Lambda_{v}\mathbf{u}$
    for all $v\in W^{1-\frac{1}{q},q}(\Gamma)$.
  5. 5.
    We finally remark that this pairing is, indeed, linear in $v$ since
    $\begin{split}\langle\operatorname{\mathfrak{tr}}_{1}(\mathbf{u}),a_{1}v_{1}+a_{2}v_{2}\rangle&=\lim\limits_{n\to\infty}\langle\operatorname{\mathfrak{tr}}_{1}(\mathbf{u}_{n}),a_{1}v_{1}+a_{2}v_{2}\rangle\\ &=\lim\limits_{n\to\infty}(a_{1}\,\langle\operatorname{\mathfrak{tr}}_{1}(\mathbf{u}_{n}),v_{1}\rangle+a_{2}\,\langle\operatorname{\mathfrak{tr}}_{1}(\mathbf{u}_{n}),v_{2}\rangle)\\ &=a_{1}\,\langle\operatorname{\mathfrak{tr}}_{1}(\mathbf{u}),v_{1}\rangle+a_{2}\,\langle\operatorname{\mathfrak{tr}}_{1}(\mathbf{u}),v_{2}\rangle\end{split}$
    for some sequence $\{\mathbf{u}_{n}\}_{n=1}^{\infty}\subset\left(\mathfrak{D}(\mkern 1.0mu\overline{\mkern-1.0mu\Omega\mkern-1.0mu}\mkern 1.0mu)\right)^{d}$ such that $\|\mathbf{u}-\mathbf{u}_{n}\|_{W^{p}(\operatorname*{div};\Omega)}\to 0$ as $n\to\infty$. Additionally,
    $\|\operatorname{\mathfrak{tr}}_{1}(\mathbf{u})\|_{W^{-1+\frac{1}{q},p}(\Gamma)}=\sup_{\|v\|\leq 1}\langle\operatorname{\mathfrak{tr}}_{1}(\mathbf{u}),v\rangle\leq C\|\mathbf{u}\|_{W^{p}(\operatorname*{div};\Omega)}$
    for some constant $C>0$ independent of $\mathbf{u}$.
So, indeed, $\operatorname{\mathfrak{tr}}_{1}(\mathbf{u})\in W^{-1+\frac{1}{q},p}(\Gamma)$ and this concludes the proof. ∎
Funding:
The research reported in this paper was funded through King Abdullah University of Science and Technology (KAUST) baseline funds and KAUST OSR-CRG2021-4674. A. M. Alharbi was additionally supported by the Islamic University of Madinah, AD no.133212662/102931.